### v1

# ELFATT: Efficient Linear Fast Attention for Vision Transformers

Abstract
The attention mechanism is the key to the success of transformers in different machine learning tasks. However, the quadratic complexity with respect to the sequence length of the vanilla softmax-based attention mechanism becomes the major bottleneck for the application of long sequence tasks, such as vision tasks. Although various efficient linear attention mechanisms have been proposed, they need to sacrifice performance to achieve high efficiency. What’s more, memory-efficient methods, such as FlashAttention-1-3, still have quadratic computation complexity which can be further improved. In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism to achieve low memory input/output operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT is FlashAttention friendly. Using FlashAttention-2 acceleration, ELFATT still offers 2-3x speedups over the vanilla softmax-based attention mechanism on high-resolution vision tasks without losing performance. Even on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W. The code of ELFATT is available at [https://github.com/Alicewithrabbit/ELFATT].
Index Terms:
Efficient attention mechanism, linear approximation, vision transformer.
I Introduction
Transformers have achieved great success in large language models (ChatGPT [1] and Llama [2]) and large vision models (SAM [3] and Sora [4]). The core technique of the transformer, the vanilla softmax-based attention mechanism, is capable of capturing the relationship between any two tokens [5]. In complex tasks, the sequence length becomes longer and longer, usually much longer than the embedding dimension. The softmax operation after the multiplication of query and key matrices makes the vanilla softmax-based attention mechanism quadratic computational complexity with respect to the sequence length. It has become a main obstacle to computational efficiency.
Acceleration methods to speed up the attention computation can be classified into two categories: (1) memory-efficient methods; (2) computation-efficient methods. Memory-efficient methods focus on optimizing memory input/output (I/O) operations to achieve almost linear complexity [6, 7, 8, 9]. FlashAttention [6, 7, 8] and FlashSigmoid [9] are representatives of memory-efficient methods. Computation-efficient methods focus on minimizing the computation bound of the attention computation by using linear approximations based on different kernel methods [10, 11, 12, 13, 14, 15, 16, 17, 18, 19], low-rank decomposition [20, 21, 15, 5], and sparse computation [22, 23, 24, 25, 26, 27, 28, 29, 30]. However, memory-efficient methods still have quadratic computational complexity, and computation-efficient methods usually have lower performance compared to the vanilla softmax-based attention mechanism.
In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to the vanilla softmax-based attention mechanism. The core idea of ELFATT is the combination of sparse computation with a linear approximation. Each ELFATT block has two parallel attention heads. One head is used to compute sparse blockify attention to introduce inductive biases, and the other one head is used to compute global linear attention to capture long-range dependencies. Both heads have almost linear complexity, and the sparse blockify attention process can be further speeded up by using the FlashAttention [6, 7, 8] mechanisms to reduce memory I/O operations. ELFATT is evaluated on different vision tasks: image classification, semantic segmentation, and object detection. What’s more, ELFATT is also evaluated on edge GPUs. Compared to state-of-the-art memory-efficient acceleration methods and computation-efficient acceleration methods, ELFATT inherits advantages of both two kinds of methods: noninferior performance compared to the vanilla softmax-based attention mechanism, low memory I/O operations, and linear computational complexity. In summary, this paper has the following contributions.
  • (i)
    A novel efficient linear fast attention (ELFATT) mechanism is proposed that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to the vanilla softmax-based attention mechanism.
  • (ii)
    The relationship between the proposed ELFATT mechanism and the vanilla softmax-based attention mechanism is analyzed and given.
  • (iii)
    An upper bound is given for the use of the proposed ELFATT mechanism to approximate the vanilla softmax-based attention mechanism.
  • (iv)
    ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism without using FlashAttention-2 and 2-3x speedups over the vanilla softmax-based attention mechanism using FlashAttention-2 on high-resolution vision tasks.
  • (v)
    ELFATT offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W on edge GPUs.
II Related Work
Since the vanilla softmax-based attention mechanism has quadratic complexity in both time and memory, depending on whether the focus is on optimizing memory complexity or time complexity, attention acceleration methods can be categorized into the following two types: (A) memory-efficient attention acceleration methods; (B) computation-efficient attention acceleration methods.
II-A Memory-Efficient Attention Acceleration Methods
Memory-efficient attention acceleration methods focus on optimizing quadratic memory complexity to almost linear complexity. One of the most representative methods is FlashAttention [6], which minimizes memory I/O access for the softmax computation of the product of query and key matrices to improve the utilization rate of GPUs to achieve almost linear memory complexity [6, 5]. FlashAttention-2 further reduces the number of floating point operations (FLOPs) of non-matrix multiplication, parallelizes both forward and backward processes according to the sequence length, and reduces the I/O access of shared memory [7]. FlashAttention-3 is proposed to fully utilize the performance of new Hopper GPUs by introducing warp-specialization, interleave block-wise matrix multiplication and softmax operations, and the support of FP8 precision [8]. FlashSigmoid investigates the feasibility of using the sigmoid function to replace the softmax function in attention computation, proposes a new regularity method for sigmoid attention to stabilize training, and introduces a memory-efficient version based on FlashAttention-2[9]. However, these memory-efficient attention acceleration methods are usually hard-aware and can not support all kinds of GPUs and their computation complexity is still quadratic.
II-B Computation-Efficient Attention Acceleration Methods
Different to hard-aware memory-efficient attention acceleration methods, computation-efficient attention acceleration methods focus on optimizing computational complexity by linear approximations and usually can be categorized into three classes: (a) kernel methods; (b) low-rank decomposition based methods; (c) sparse methods.
Kernel methods usually change the order of nonlinear normalization and multiplication of the query matrix and key matrix to achieve linear computational complexity. FLatten [10] introduces a cubic rectified linear unit (ReLU) [31] based feature map and uses it to perform nonlinear normalization on the query matrix and key matrix, respectively, before the attention computation. The linear transformer [13] introduces a feature map based on the exponential linear unit (elu) [32] activation function. cosFormer [17] introduces a cos-based nonlinear feature map to obtain a linear approximation of the vanilla softmax-based attention mechanism. Efficient attention [18] performs softmax normalization on the query matrix and the key matrix, respectively, to achieve linear complexity [18]. However, this linearization introduces noise from two aspects: (1) Separated softmax normalization will reduce the similarity of elements in the query matrix and key matrix that are both negative at the same position; (2) Large negative values are suppressed to small positive values. The above two types of noise result in decreased discrimination between tokens and incorrect concentration of attention. Agent [19] introduces a small agent matrix, which is obtained by performing the pooling operation on the query matrix. This agent matrix is used as the auxiliary key matrix to be multiplied with the query matrix to reduce its dimenison before the softmax normalization. Similarly, it is also used to reduce the dimension of the key matrix before its softmax normalization. This agent matrix served as a bridge between two independent softmax normalization processes to reduce noise.
Low-rank decomposition based methods usually take the softmax normalization of the product of query and key matrices as a whole for decomposition. Nyströmformer [21] introduces Nyström approximation to perform low-rank decomposition of the softmax normalization of the product of query and key matrices. For calculating an attention score matrix, it needs to perform an iterative inverse approximation and operate softmax normalization three times (each softmax normalization still involves a product of two matrices), which makes its acceleration effect not significant. Skeleton decomposition-based self-attention uses a simplified inverse approximation method based on the permuted diagonal matrix [20]. Interactive multihead self-attention (iMHSA) introduces several linear layers to perform head interactions instead of performing an iterative inverse approximation [33]. CUR decomposition based self-attention (CURSA) [5] introduces the CUR decomposition to replace the Nyström approximation and reduces the number of matrix multiplication.
Sparse methods usually separate the original sequences into smaller blocks using sliding windows and perform attention computation within these blocks to reduce complexity. Each token in a sequence is only affected by several tokens, not all tokens; hence, this kind of sparse methods are also called local attention mechanism. Swin [22] and Longformer [26] are pioneers in introducing sliding windows into attention computation for vision tasks and language tasks, respectively. To address the information loss introduced by the local windows, Swin introduces a shifted window mechanism to capture cross-block information, while Longformer selects some tokens as global tokens, which have effects on all tokens, to compensate for the global information loss. NesT [34] further simplifies the shifted window mechanism through simple spatial operations. CSWin [23] introduces a cross-shaped window mechanism for 2D sequences. The cross-shaped window mechanism separates a 2D sequence into horizontal and vertical stripes, respectively, and performs parallel attention computation within these stripes in horizontal and vertical directions, which can simultaneously obtain local inductive biases and cross-block (global) information. In addition to using the deterministic global token selection mechanism of Longformer, Big Bird [29] also randomly selects some global tokens to enhance performance. Similarly, the sparse transformer [27] proposed by OpenAI also introduces several fixed-step tokens, which is similar to dilated convolution [35] in convolutional neural networks (CNNs) to capture long-range information. Reformer [30] finds the nearest neighbors for each token to calculate local attention scores using the locality-sensitive-hashing (LSH) algorithm.
III Methods
III-A Vanilla Softmax-Based Attention Mechanism
For clarity, we assume that all vectors appear in this paper are row vectors. For any two same sized tokens $\textit{{x}}\in\mathbb{R}^{c}$ and $\textit{{y}}\in\mathbb{R}^{c}$, their attention similarity score can be calculated as follows,
$a={\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right),$ (1)
where ${\rm exp}\left(\cdot\right)$ is an element-wise exponential function and is conducted after the inner product of two tokens. For given two $m$-tuples $\{\textit{{x}}_{1},\dots,\textit{{x}}_{m}\}$ and $\{\textit{{y}}_{1},\dots,\textit{{y}}_{m}\}$ with $\textit{{x}}_{i},\textit{{y}}_{i}\in\mathbb{R}^{c}$ and $i=1,2,\dots,m$, we need to calculate $m^{2}$ pairs of attention similarity scores. Currently, a long sequence length $m$ is preferred in large models and $m\gg c$, therefore, the computational complexity of the vanilla softmax-based attention mechanism is quadratic with respect to the sequence length $m$.
III-B General Attention Mechanism
Eq. (1) can be rewritten as a more general form [13] as follows,
$a={\rm sim}\left(\textit{{x}},\textit{{y}}\right),$ (2)
where ${\rm sim}(\cdot)$ is a non-negative similarity function and it satisfies the definition of a kernel function $G(\textit{{x}},\textit{{y}}):\mathbb{R}^{c}\times\mathbb{R}^{c}\rightarrow\mathbb{R}_{+}$ with +=$\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$.
III-C Kernelized Attention Mechanism
If we have such a kernel with a non-negative feature map $\phi$, Eq. (2) can be written as follows,
$a=\phi\left(\textit{{x}}\right)\phi\left(\textit{{y}}\right)^{\top}.$ (3)
Performer [11] has proved that one of the best choices of the non-negative feature map $\phi$ for Eq. (3) is $\rm exp\left(\cdot\right)$. Performer obtains an exact alternative of Eq. (1) using random feature maps as follows,
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)=\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}\right)\right],$ (4)
where $e\left(\textit{{x}}\right)={\rm exp}\left(-\frac{\|\textit{{x}}\|^{2}}{2}\right)$, $e\left(\textit{{y}}\right)={\rm exp}\left(-\frac{\|\textit{{y}}\|^{2}}{2}\right)$, $\textbf{0}_{c}\in\mathbb{R}^{c}$ is the zero vector, and $\textbf{I}_{c}\in\mathbb{R}^{c\times c}$ is the identity matrix. If the attention scores for all pairs of x and $\textit{{y}}_{i}$ ($i=1,2,\dots,m$) using Eq. (4) are obtained, after the following normalization $e\left(\textit{{x}}\right)$ will be canceled.
$\begin{split}&\frac{{\rm exp}\left(\textit{{x}}\textit{{y}}_{i}^{\top}\right)}{\sum_{j=1}^{m}{{\rm exp}\left(\textit{{x}}\textit{{y}}_{j}^{\top}\right)}}\\ &=\frac{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{i}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}_{i}\right)\right]}{\sum_{j=1}^{m}{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{j}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}_{j}\right)\right]}}\\ &=\frac{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{i}^{\top}\right)e\left(\textit{{y}}_{i}\right)\right]}{\sum_{j=1}^{m}{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{j}^{\top}\right)e\left(\textit{{y}}_{j}\right)\right]}}.\end{split}$
Hence, $e\left(\textit{{x}}\right)$ has no effect on the final attention score. Eq. (4) can be written as follows,
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)=\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{y}}\right)\right].$ (5)
Eq. (5) only holds when taking the sum of an infinite number of random vectors $\omega$. To avoid performing summation of infinite terms, Performer samples $c\times\log(c)$ random vectors $\omega$ to ensure a low approximation error. If $e\left(\textit{{y}}\right)$ is much smaller than ${\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)$, Eq. (5) can be further simplified and approximated as follows,
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)\right].$ (6)
Efficient attention (EFFATT) [18] is a special case of Eq. (6). It simplifies $\omega$ to a one-hot vector and only uses $c$ one-hot vectors in Eq. (6) to obtain an approximation as follows,
${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\frac{1}{c}{\rm exp}\left(\textit{{x}}\right){\rm exp}\left(\textit{{y}}\right)^{\top}.$ (7)
Its approximation error was studied and obtained in [5]. Eq. (7) has a problem of concentration reduction of attention maps [5] and Performer also has this problem when the number of $\omega$ sampled is too small. When the number of $\omega$ sampled is too large, Performer may be slower than the vanilla softmax-based attention mechanism.
III-D Efficient Linear Fast Attention Mechanism
To address the problem of concentration reduction of Eq. (7), in this paper, we propose a novel attention mechanism as follows,
$\begin{split}{\rm exp}&\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$ (8)
where $\textit{{Q}}=[\bar{\textit{{Q}}},\tilde{\textit{{Q}}}]\in\mathbb{R}^{m\times c}$, $\textit{{K}}=[\bar{\textit{{K}}},\tilde{\textit{{K}}}]\in\mathbb{R}^{m\times c}$, $\textit{{V}}=[\bar{\textit{{V}}},\tilde{\textit{{V}}}]\in\mathbb{R}^{m\times c}$, $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{Q}}}\in\mathbb{R}^{m\times c_{2}}$, $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{K}}}\in\mathbb{R}^{m\times c_{2}}$, $\bar{\textit{{V}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{V}}}\in\mathbb{R}^{m\times c_{2}}$, $c=c_{1}+c_{2}$, $f\left(\cdot\right)$ is a blockify function to separate a matrix with a size of $m\times c_{2}$ into $b$ blocks (each block has a size of $(m/b)\times c_{2}$), and $g\left(\cdot\right)$ is an unblockify function to unblock $b$ blocks to a single matrix with a size of $m\times c_{2}$. Eq. (8) can be derived as follows,
$\begin{split}&{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\\ &=\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\textit{{V}}\\ &={\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\left[\bar{\textit{{V}}},\tilde{\textit{{V}}}\right]\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right],\\ \end{split}$
where $\odot$ denotes the Hadamard product [36], ${\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\tilde{\textit{{V}}}$ is equivalent to $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)$, and $\textit{{Z}}\in\mathbb{R}^{m\times m}$ is a matrix as follows,
$\textit{{Z}}=\textbf{I}_{b}\otimes\textit{{U}}_{(m/b)},$
with $\otimes$ denoting the Kronecker product [36] and $\textit{{U}}_{(m/b)}\in\mathbb{R}^{(m/b)\times(m/b)}$ being the all-ones matrix. It is obvious to see that
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\right.\\ &\quad\left.\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right],\end{split}$
which implies that
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi},\end{split}$ (9)
where $\xi=2$ denotes the spectral norm and $\xi=F$ denotes the Frobenius norm. For given two matrices $\textit{{A}}\in\mathbb{R}^{m\times n}$ and $\textit{{B}}\in\mathbb{R}^{n\times p}$, it follows from [37, Section 2.3.1] that $\|\textit{{A}}\textit{{B}}\|_{\xi}\leq\|\textit{{A}}\|_{\xi}\|\textit{{B}}\|_{\xi}$.
For the second term in the right-hand side of Inequality (9), it is easy to see that
$\begin{split}&\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
For the first term in the right-hand side of Inequality (9), we have the following theorem.
Theorem 1.
Let $\emph{{{U}}}_{m}\in\mathbb{R}^{m\times m}$ be an all-ones matrix. For any two vectors $\textit{{q}}\in\mathbb{R}^{c}$ from $\textit{{Q}}\in\mathbb{R}^{m\times c}$ and $\textit{{k}}\in\mathbb{R}^{c}$ from $\textit{{K}}\in\mathbb{R}^{m\times c}$, let $\mathfrak{M}>0$ and $\mathscr{M}>0$ be the maximum and minimum of ${\rm exp}\left(\textit{{q}}\textit{{k}}^{\top}+0.5-\left(q_{i}+k_{i}\right)\right)$, respectively, and $q_{i}$ and $k_{i}$ are the elements at position $i$ of vectors q and k, respectively. If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
Before proving Theorem 1, we introduce the following lemma (see [5]).
Lemma 1.
For any two vectors $\textit{{q}}\in\mathbb{R}^{c}$ from $\textit{{Q}}\in\mathbb{R}^{m\times c}$ and $\textit{{k}}\in\mathbb{R}^{c}$ from $\textit{{K}}\in\mathbb{R}^{m\times c}$, let
$\begin{split}D_{\textit{{q}},\textit{{k}}}&=\max_{i=1,2,\dots,c}{\rm exp}(\textit{{q}}\textit{{k}}^{\top}+0.5-\left(q_{i}+k_{i}\right))>0,\\ d_{\textit{{q}},\textit{{k}}}&=\min_{i=1,2,\dots,c}{\rm exp}(\textit{{q}}\textit{{k}}^{\top}+0.5-\left(q_{i}+k_{i}\right))>0,\end{split}$
where $q_{i}$ and $k_{i}$ are the elements at position $i$ of vectors q and k, respectively. The following inequalities hold,
$\begin{split}\frac{{\rm exp}\left(\textit{{q}}\right){\rm exp}\left(\textit{{k}}\right)^{\top}}{{\rm exp}\left(\textit{{q}}\textit{{k}}^{\top}\right)}&\leq\frac{c}{d_{\textit{{q}},\textit{{k}}}{\rm exp}\left(-0.5\right)},\\ \frac{{\rm exp}\left(\textit{{q}}\right){\rm exp}\left(\textit{{k}}\right)^{\top}}{{\rm exp}\left(\textit{{q}}\textit{{k}}^{\top}\right)}&\geq\frac{c}{D_{\textit{{q}},\textit{{k}}}{\rm exp}\left(-0.5\right)}.\end{split}$
The following corollary is easily obtained from Lemma 1.
Corollary 1.
For two matrices $\textit{{Q}}\in\mathbb{R}^{m\times c}$ and $\textit{{K}}\in\mathbb{R}^{m\times c}$, let
$\begin{split}\mathfrak{M}&=\max_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c}{\rm exp}(\textit{{q}}_{i_{1}}\textit{{k}}_{i_{2}}^{\top}+0.5-\left(q_{i_{1}j}+k_{i_{2}j}\right))>0,\\ \mathscr{M}&=\min_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c}{\rm exp}(\textit{{q}}_{i_{1}}\textit{{k}}_{i_{2}}^{\top}+0.5-\left(q_{i_{1}j}+k_{i_{2}j}\right))>0,\end{split}$
where $q_{i_{1}j}$ and $k_{i_{2}j}$ are the elements at position $j$ of the vector $\textit{{q}}_{i_{1}}=\textit{{Q}}(i_{1},:)$ and the vector $\textit{{k}}_{i_{2}}=\textit{{K}}(i_{2},:)$, respectively. The following inequalities hold,
$\begin{split}\frac{\left\|{\rm exp}\left(\textit{{Q}}\right){\rm exp}\left(\textit{{K}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\right\|_{\xi}}&\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)},\\ \frac{\left\|{\rm exp}\left(\textit{{Q}}\right){\rm exp}\left(\textit{{K}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\right\|_{\xi}}&\geq\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}.\end{split}$
Proof.
For any matrices $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{Q}}}\in\mathbb{R}^{m\times c_{2}}$, $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, and $\tilde{\textit{{K}}}\in\mathbb{R}^{m\times c_{2}}$, we have
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &=\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)+\right.\\ &\left.{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\\ &\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\right\|_{\xi}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
Following from Corollary 1, if $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
The proof is completed. ∎
We now consider the upper bound for the right-hand side of Inequality (9). If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\right.\\ &\left.\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (10)
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\right.\\ &\left.\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (11)
Eq. (8) denotes the single-head vanilla softmax-based attention mechanism approximated by using ELFATT, and Inequality (9) denotes the corresponding approximation error. From the right-hand side of Eq. (8), each ELFATT attention process consists of two heads. Hence, ELFATT can directly become an approximation of the double-head vanilla softmax-based attention mechanism as follows,
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right].\end{split}$ (12)
It is obvious to see that
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\\ &\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\right.\\ &\left.\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right].\end{split}$
Hence, the corresponding approximation error is bounded as follows,
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\right.\\ &\left.\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (13)
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\right.\\ &\left.\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\leq\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-\emph{{{U}}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (14)
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\right.\\ &\left.\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\leq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-\emph{{{U}}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (15)
Inequalities (13-15) give a tighter bound than Inequalities (9-11).
Since each ELFATT attention block corresponds to two parallel heads, $s$ ELFATT will be needed for the approximation of $2\times s$ heads of vanilla softmax-based attention according to Eq. (12). For the approximation of $12\times s+$ 1 heads of vanilla softmax-based attention, $12\times s+$ 1 ELFATT will be needed according to Eq. (8).
III-E Positional Encoding
Different positional encoding mechanisms such as absolute positional encoding [38], relative positional encoding [39], and conditional positional encoding [40], have been proposed to make use of the ordering information of sequences in vision transformers [38, 23, 40]. Among these positional encoding mechanisms, the locally enhanced positional encoding (LePE) [23] mechanism shows more powerful local positional information enhancement and brings a higher performance gain for vision transformers. Hence, LePE is selected as the positional encoding mechanism for ELFATT. After introducing LePE, Eq. (8) will become as follows,
$\begin{split}&{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}+L\left(\textit{{V}}\right)\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)\right],\end{split}$ (16)
(a) Vanilla attention mechanism
(b) ELFATT
Figure 1: The comparison of the vanilla attention mechanism and ELFATT. For an input embedding matrix $\textit{{H}}\in\mathbb{R}^{m\times n}$, each ELFATT block will process it in two parallel attention heads to obtain two new embedding matrices, $\bar{\textit{{H}}}\in\mathbb{R}^{m\times c_{1}}$ and $\tilde{\textit{{H}}}\in\mathbb{R}^{m\times c_{2}}$, respectively. After a concatenation operation, the ELFATT block will output the updated embedding matrix $\textit{{H}}=[\bar{\textit{{H}}},\tilde{\textit{{H}}}]\in\mathbb{R}^{m\times c}$. Note: ${\rm cat}$ denotes a concatenation operation for concatenating two matrices on the column dimension into a single matrix.
Figure 2: Some visual comparison of class activation map (CAM) based attention results of different attention mechanisms obatined using Score-CAM [41]. Note: Backbone used is CSWin-T24181 [10].
Figure 3: More visual comparison of class activation map (CAM) based attention results of different attention mechanisms obatined using Score-CAM [41]. Note: Backbone used is CSWin-T24181 [10].
Eq. (12) will become as follows,
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}+L\left(\tilde{\textit{{V}}}\right)\right]\approx\\ &\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right.\right.\\ &\left.\left.+L\left(f(\tilde{\textit{{V}}})\right)\right)\right],\end{split}$ (17)
where $L(\cdot)$ denotes a depthwise convolution operation with a kernel size of 3. Fig. 1 shows the comparison of the vanilla attention mechanism and ELFATT. For an input embedding matrix $\textit{{H}}\in\mathbb{R}^{m\times n}$, each ELFATT block will process it in two parallel attention heads to obtain two new embedding matrices, $\bar{\textit{{H}}}\in\mathbb{R}^{m\times c_{1}}$ and $\tilde{\textit{{H}}}\in\mathbb{R}^{m\times c_{2}}$, respectively. After a concatenation operation, the ELFATT block will output the updated embedding matrix $\textit{{H}}=[\bar{\textit{{H}}},\tilde{\textit{{H}}}]\in\mathbb{R}^{m\times c}$.
III-F Complexity Analysis
According to Eqs. (16) and (17), the complexity of ELFATT can be divided into two parts: (a) the global linear attention head ${\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right)$; (b) the local blockify attention head $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)$. The complexity of the global linear attention head is $O(m\times c_{1}^{2})$ in the case of $m>c_{1}$, and $O(m^{2}\times c_{1})$ in the case of $m\leq c_{1}$. The complexity of the local blockify attention head is $O((m^{2}/b)\times c_{2})$. If $m/b\ll m$, the local blockify attention head will also achieve almost linear complexity.
IV Experiments and Results
We evaluated ELFATT in commonly used vision tasks: image classification (ImageNet-1K [42]), semantic segmentation (ADE20K [43]), and object detection (MS COCO 2017 [44]). We compared ELFATT with the vanilla softmax-based attention mechanism [38], the memory-efficient attention mechanism (FlashAttention-2 [7]), local window-based attention mechanisms (Swin [22] and CSWin [23]), and kernel-based attention mechanisms (Agent [19], EFFATT [18], and FLatten [10]). The backbone ViT architectures used to evaluate the different attention mechanisms are: Swin-T [22], Swin-B [22], CSWin-T24181 [10], and CSWin-B36292 [10]. The original Swin-T, CSWin-T24181, and CSWin-B36292 are called Swin-T-LOCAL, Swin-B-LOCAL, CSWin-T24181-LOCAL, and CSWin-B36292-LOCAL, respectively. The backbones after replacing local window-based attention mechanisms with the vanilla softmax-based attention mechanism [38] are called Swin-T-GLOBAL, Swin-B-GLOBAL, CSWin-T24181-GLOBAL, and CSWin-B36292-GLOBAL, respectively. The backbones after replacing local window-based attention mechanisms with kernel-based attention mechanisms (Agent [19], EFFATT [18], and FLatten [10]) are called Swin-T-Agent, Swin-T-EFFATT, Swin-T-FLatten, Swin-B-Agent, Swin-B-EFFATT, Swin-B-FLatten, CSWin-T24181-Agent, CSWin-T24181-EFFATT, CSWin-T24181-FLatten, CSWin-B36292-Agent, CSWin-B36292-EFFATT, and CSWin-B36292-FLatten, respectively. The backbones after replacing local window-based attention mechanisms with ELFATT are called Swin-T-ELFATT, Swin-B-ELFATT, CSWin-T24181-ELFATT, and CSWin-B36292-ELFATT, respectively. For the ImageNet-1K image classification task, we used the same training settings and data augmentation methods from [23] to train all models from scratch using mixed precision. For the ADE20K semantic segmentation task [43] and the MS COCO 2017 object detection task [44], we used the same training settings and data augmentation methods from [45] to fine-tune the weights of all models obtained from the ImageNet-1K image classification task using mixed precision. We also compared ELFATT with ConvNeXt-T [46] and VMamba-T [45] on ImageNet-1K, ADE20K, and MS COCO 2017. The experiments of ImageNet-1K, ADE20K, and MS COCO 2017 were carried out on 8 NVIDIA vGPU (32 GB) GPUs. Inference throughput comparison experiments were carried out on 1 NVIDIA H20 (96 GB) GPU. Besides high-performance computing applications, we also compared ELFATT with state-of-the-art attention mechanisms on edge GPUs (NVIDIA Jetson AGX Orin/NVIDIA Jetson Nano GPU). We used FlashAttention-2 [7] to speed up all models that are compatible with FlashAttention. The PyTorch implementation of ELFATT is available at [https://github.com/Alicewithrabbit/ELFATT].
IV-A Image Classification Performance
Table I shows the performance comparison of different attention mechanisms on ImageNet-1K. From Table I, the vanilla softmax-based attention mechanism (Swin-T-GLOBAL and CSWin-T24181-GLOBAL) can outperform most linear attention mechanisms using the same architecture. Only Agent and ELFATT achieve the most close performance compared to the vanilla softmax-based attention mechanism in this paper. As to the inference speed comparison, the proposed ELFATT achieves the highest inference throughput (frame per second, FPS) than all other attention mechanisms when using CSWin-T24181 as the backbone. ELFATT achieves almost the same speed as EFFATT and is significantly faster than other attention mechanisms when using Swin-T as the backbone. ELFATT offers almost a 2x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2. The vanilla softmax-based attention mechanism using FlashAttention-2 is still 0.1-0.2x slower than ELFATT without using FlashAttention-2. With the use of FlashAttention-2 to optimize memory operations, ELFATT can be further accelerated and offers 1.2-1.3x speedups over the vanilla softmax-based attention mechanism. Figs. 2-3 show the visual comparison of different attention mechanisms using CSWin-T24181 as the backbone. ELFATT shows a much closer attention map to the vanilla softmax-based attention mechanism than other attention mechanisms. For the larger backbone, CSWin-B36292, ELFATT still achieves state-of-the-art (SOTA) performance. ELFATT offers a 2.1/1.4x speedup over the vanilla softmax-based attention mechanism with/without using FlashAttention-2 for an input resolution of $224^{2}$ and a 3.3/1.8x speedup over the vanilla softmax-based attention mechanism with/without using FlashAttention-2 for an input resolution of $384^{2}$. In addition, ELFATT achieves significant leading performance under the backbone of Swin-B, offering a 1.8x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 for an input resolution of $224^{2}$ and a 2.8x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 for an input resolution of $384^{2}$. After using FlashAttention-2, it matches the speed of EFFATT which is a real linear attention mechanism. Also, as shown in Table I, it can be seen that although some methods have lower FLOPs than ELFATT, their actual speed is slower than ELFATT which is consistent with the observations and conclusions of [47].
TABLE I: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of different attention mechanisms on ImageNet-1K. Note: Inference throughput of tiny models is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU and inference throughput of base models is obtained using a batch size of 256 for $224^{2}$ resolution and 32 for $384^{2}$ resolution with mixed precision on a single NVIDIA H20 (96 GB) GPU. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-B36292-Agent $224^{2}$ 84.7 930/994 imgs/s 73M 14.92G/14.49G
CSWin-B36292-EFFATT $224^{2}$ 84.4 985/1059 imgs/s 73M 14.98G/14.53G
CSWin-B36292-ELFATT $224^{2}$ 84.7 1000/1187 imgs/s 73M 15.47G/14.46G
CSWin-B36292-FLatten $224^{2}$ 84.5 814/864 imgs/s 75M 14.96G/14.52G
CSWin-B36292-GLOBAL $224^{2}$ 84.7 478/879 imgs/s 73M 22.33G/14.39G
CSWin-B36292-LOCAL $224^{2}$ 84.4 941/1037 imgs/s 73M 15.03G/14.39G
CSWin-B36292-Agent $384^{2}$ 85.8 246/276 imgs/s 73M 46.33G/42.57G
CSWin-B36292-EFFATT $384^{2}$ 85.7 294/331 imgs/s 73M 46.57G/42.69G
CSWin-B36292-ELFATT $384^{2}$ 85.8 272/355 imgs/s 73M 51.21G/42.48G
CSWin-B36292-FLatten $384^{2}$ 85.5 238/266 imgs/s 78M 46.43G/42.67G
CSWin-B36292-GLOBAL $384^{2}$ 85.9 82/201 imgs/s 73M 110.89G/42.28G
CSWin-B36292-LOCAL $384^{2}$ 85.5 276/323 imgs/s 73M 47.06G/42.28G
CSWin-T24181-Agent $224^{2}$ 83.1 2297/2425 imgs/s 20M 4.31G/4.14G
CSWin-T24181-EFFATT $224^{2}$ 82.6 2394/2526 imgs/s 20M 4.35G/4.17G
CSWin-T24181-ELFATT $224^{2}$ 83.1 2603/2856 imgs/s 20M 4.44G/4.13G
CSWin-T24181-FLatten $224^{2}$ 83.1 1934/2025 imgs/s 21M 4.34G/4.16G
CSWin-T24181-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
CSWin-T24181-LOCAL $224^{2}$ 82.7 2330/2519 imgs/s 20M 4.36G/4.09G
Swin-B-Agent $224^{2}$ 84.0 1367/ —   imgs/s 88M 15.44G/   —
Swin-B-EFFATT $224^{2}$ 84.2 1439/1536 imgs/s 88M 15.69G/15.33G
Swin-B-ELFATT $224^{2}$ 84.5 1314/1497 imgs/s 91M 16.46G/15.68G
Swin-B-FLatten $224^{2}$ 83.8 1226/ —   imgs/s 89M 15.46G/   —
Swin-B-GLOBAL $224^{2}$ 84.2 743/1325 imgs/s 88M 21.57G/15.19G
Swin-B-LOCAL $224^{2}$ 83.5 1351/ —   imgs/s 88M 15.47G/   —
Swin-B-Agent $384^{2}$ 84.9 372/ — imgs/s 88M 46.34G/   —
Swin-B-EFFATT $384^{2}$ 85.3 433/481 img/s 88M 48.18G/45.04G
Swin-B-ELFATT $384^{2}$ 85.5 372/457 img/s 91M 52.79G/46.08G
Swin-B-FLatten $384^{2}$ 85.0 353/ — imgs/s 91M 46.49G/   —
Swin-B-GLOBAL $384^{2}$ 85.3 134/313 img/s 88M 99.76G/44.64G
Swin-B-LOCAL $384^{2}$ 84.5 357/ — imgs/s 88M 47.19G/   —
Swin-T-Agent $224^{2}$ 82.6 2847/ —   imgs/s 29M 4.53G/ —
Swin-T-EFFATT $224^{2}$ 82.1 3165/3282 imgs/s 28M 4.55G/4.45G
Swin-T-ELFATT $224^{2}$ 82.7 2884/3159 imgs/s 30M 4.99G/4.67G
Swin-T-FLatten $224^{2}$ 82.1 2502/ —   imgs/s 29M 4.50G/ —
Swin-T-GLOBAL $224^{2}$ 82.4 1269/2571 imgs/s 28M 8.81G/4.38G
Swin-T-LOCAL $224^{2}$ 81.4 2943/ —   imgs/s 28M 4.51G/ —
ConvNeXt-T $224^{2}$ 82.1 3911/ —   imgs/s 29M 4.47G/ —
VMamba-T $224^{2}$ 82.5 1837/ —   imgs/s 30M 4.84G/ —
TABLE II: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT using different block sizes in each level of CSWin-T24181 on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Block size of each level Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
Level 1 Level 2 Level 3 Level 4
49 49 49 49 $224^{2}$ 82.5 2619/2805 imgs/s 20M 4.37G/4.13G
49 49 196 49 $224^{2}$ 83.1 2603/2856 imgs/s 20M 4.44G/4.13G
49 196 196 49 $224^{2}$ 82.8 2552/2863 imgs/s 20M 4.50G/4.13G
196 196 196 49 $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
IV-B Ablation Study of Block Size
To validate the effect of the block size on the performance of ELFATT, we performed an ablation study using different block sizes at each level of CSWin-T24181-ELFATT on ImageNet-1K. As shown in Table II, with the increasing number of block sizes, the inference speed of ELFATT will slow down, while the classification accuracy of ELFATT shows an increasing trend. ELFATT using 49-49-196-49 and 196-196-196-49 blocks achieve the best performance in terms of speed and accuracy. It seems that FlashAttention-2 acceleration is more efficient for some block sizes (with FlashAttention-2 acceleration, the block size of 196 is faster than the block size of 49 which may be caused by GPU architectures). Hence, the block sizes for ELFATT can be determined according to the performance and efficiency requirements.
IV-C Ablation Study of LePE
To validate the effect of LePE on performance, we compared the performance of ELFATT and the vanilla softmax-based attention mechanism using or without using LePE in CSWin-T24181 on ImageNet-1K. As shown in Table III, using LePE, the performance of the vanilla softmax-based attention mechanism improves from 82.9 to 83.1 and the performance of ELFATT also improves from 82.9 to 83.1, respectively, further demonstrating the powerful local positional information enhancement of LePE. More details about LePE can be found in [23].
TABLE III: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT and the vanilla softmax-based attention mechanism using or without using LePE in CSWin-T24181 on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “w/o” denotes “without”, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-T24181-ELFATT $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
CSWin-T24181-ELFATT-w/o-LePE $224^{2}$ 82.9 2801/3271 imgs/s 20M 4.54G/4.12G
CSWin-T24181-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
CSWin-T24181-GLOBAL-w/o-LePE $224^{2}$ 82.9 1332/2296 imgs/s 20M 7.58G/4.08G
TABLE IV: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by using the ELFATT modules to replace the vanilla softmax-based attention modules in different levels of CSWin-T24181-GLOBAL on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration. ✓denotes the level consists of full ELFATT modules and ✔ denotes half of this level is composed of ELFATT modules and the other half is composed of vanilla softmax-based attention (VaniATT) modules. The order of composition is: VaniATT-ELFATT-...-VaniATT-ELFATT.
Levels with using the ELFATT module Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
Level 1 Level 2 Level 3 Level 4
$224^{2}$ 83.0 2257/2865 imgs/s 20M 5.17G/4.10G
$224^{2}$ 83.0 2467/2918 imgs/s 20M 4.63G/4.12G
$224^{2}$ 82.6 2555/2842 imgs/s 20M 4.48G/4.15G
$224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
$224^{2}$ 82.6 2552/2835 imgs/s 20M 4.48G/4.15G
CSWin-T24181-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
TABLE V: The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by CSWin-T24181-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Combinations of $c_{1}$ and $c_{2}$ Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
$c_{1}$ $c_{2}$
$0\times c$ $1\times c$ $224^{2}$ 82.7 2331/2880 imgs/s 20M 4.76G/4.09G
0.25×0.25$\times$0.25 × c $0.75\times c$ $224^{2}$ 83.0 2530/2835 imgs/s 20M 4.65G/4.10G
0.5×0.5$\times$0.5 × c $0.5\times c$ $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
0.75×0.75$\times$0.75 × c $0.25\times c$ $224^{2}$ 83.1 2458/2796 imgs/s 20M 4.48G/4.18G
1×1$\times$1 × c $0\times c$ $224^{2}$ 82.6 2394/2526 imgs/s 20M 4.35G/4.17G
IV-D Ablation Study of Number of Levels Using the ELFATT Module
To validate the effect of the number of ELFATT modules used to replace the vanilla softmax-based attention (VaniATT) modules at different levels of vision transformers, we compared the performance of CSWin-T24181 using a different number of ELFATT modules at different levels. As shown in Table IV, with the number of levels using ELFATT modules increasing, the inference speed increases without using FlashAttention-2 acceleration. The difference between CSWin-T24181 using 3 levels of ELFATT modules and 4 levels of ELFATT modules is not significant. The fourth level of CSWin-T24181 is composed of only one module and the sequence length is only 49 which is too short to achieve a significant acceleration effect. Another thing that can be found in Table IV is that with an increase in the number of levels using ELFATT modules, CSWin-T24181 using FlashAttention-2 acceleration achieves the fastest inference speed when using two levels of ELFATT modules. Because the sequence lengths of the first two levels are 3136 and 784, respectively, which are significantly longer than the last two levels (the $3^{\rm rd}$ level: 196, and the $4^{\rm th}$ level: 49). That is also the reason why some efficient attention mechanisms, such as Agent and FLatten, only replace some of levels by their efficient attention modules. We also observed that with an increasing number of levels using ELFATT modules, the performance shows a gradual decline. To address this defect, we introduced a hybrid architecture in the third level which replaces half of the ELFATT modules at this level by vanilla softmax-based attention modules. The order of composition is as follows: VaniATT-ELFATT-...-VaniATT-ELFATT. The reason is that the third level of the vision transformer is usually much deeper than other levels and the sequence of this level is also much shorter than the first two levels. VaniATT at this level will not affect the speed too much and can help the model to converge faster. Swin-T-ELFATT and CSWin-B36292-ELFATT also use a pure ELFATT architecture in the first two levels and a hybrid architecture in the third level. All variants of CSWin-T24181 using ELFATT modules to replace VaniATT modules are faster than CSWin-T24181-GLOBAL of which all levels are composed of VaniATT modules.
TABLE VI: The comparison of object detection performance of all methods on MS COCO 2017. Note: FLOPs are calculated using an input size of $1280\times 800$. ‘1$\times$’ denotes the fine-tuning training schedule with 12 epochs and ‘3$\times$×MS’ represents fine-tuning using the multiscale training schedule with 36 epochs. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. APb denotes box average precision and APm denotes mask average precision. Inference throughput is obtained using a batch size of 1 with mixed precision on a single NVIDIA H20 (96 GB) GPU. “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
Mask-RCNN [48] 1$\times$ schedule
[…]
CSWin-T24181-Agent 46.8 68.9 51.3 42.3 65.9 45.3 19/20 imgs/s 40M 273.87G/254.51G
CSWin-T24181-EFFATT 46.1 68.3 50.5 41.9 65.5 45.3 32/36 imgs/s 40M 329.95G/255.20G
CSWin-T24181-ELFATT 47.0 69.2 51.4 42.6 66.4 45.9 30/33 imgs/s 40M 334.86G/254.40G
CSWin-T24181-FLatten 46.6 68.8 51.0 42.2 65.7 45.3 26/28 imgs/s 41M 274.41G/255.05G
CSWin-T24181-GLOBAL 47.0 69.1 51.9 42.6 66.1 45.9 7/18 imgs/s 40M 1712.76G/253.56G
CSWin-T24181-LOCAL 46.5 68.5 51.0 42.1 65.6 45.3 26/28 imgs/s 40M 281.45G/253.57G
Swin-T-Agent 44.6 67.5 48.7 40.7 64.4 43.4 5/— imgs/s 48M 278.42G/ —
Swin-T-EFFATT 44.7 67.0 48.9 41.1 64.0 44.4 40/46 imgs/s 48M 301.89G/261.95G
Swin-T-ELFATT 46.1 68.3 50.8 42.1 65.4 45.3 39/45 imgs/s 50M 311.39G/266.43G
Swin-T-FLatten 44.2 67.3 48.5 40.2 63.8 43.0 41/— imgs/s 49M 266.43G/ —
Swin-T-GLOBAL 45.4 67.9 49.7 41.6 65.0 44.8 6/17 imgs/s 48M 2106.75G/260.48G
Swin-T-LOCAL 42.7 65.2 46.8 39.3 62.2 42.2 45/— imgs/s 48M 267.01G/ —
ConvNeXt-T 44.2 66.6 48.3 40.1 63.3 42.8 44/— imgs/s 48M 262.29G/ —
VMamba-T 47.4 69.5 52.0 42.7 66.3 46.0 35/— imgs/s 50M 271.16G/ —
Mask-RCNN [48] 3$\times$×MS schedule
[…]
CSWin-T24181-Agent 49.3 70.8 53.9 43.9 67.9 47.3 19/20 imgs/s 40M 273.87G/254.51G
CSWin-T24181-EFFATT 48.5 70.0 53.2 43.4 67.3 46.9 32/36 imgs/s 40M 329.95G/255.20G
CSWin-T24181-ELFATT 49.4 70.9 54.4 44.0 68.0 47.5 30/33 imgs/s 40M 334.86G/254.40G
CSWin-T24181-FLatten 48.9 70.8 53.5 43.9 67.9 47.3 26/28 imgs/s 41M 274.41G/255.05G
CSWin-T24181-GLOBAL 48.8 70.0 53.5 43.6 67.4 47.1 7/18 imgs/s 40M 1712.76G/253.56G
CSWin-T24181-LOCAL 49.3 70.8 54.3 44.0 67.8 47.5 26/28 imgs/s 40M 281.45G/253.57G
Swin-T-Agent 47.3 69.5 51.9 42.7 66.4 46.2 5/— imgs/s 48M 278.42G/ —
Swin-T-EFFATT 47.6 69.4 52.6 42.7 65.9 46.1 40/46 imgs/s 48M 301.89G/261.95G
Swin-T-ELFATT 48.5 70.4 53.4 43.6 67.3 47.3 39/45 imgs/s 50M 311.39G/266.43G
Swin-T-FLatten 46.5 68.5 50.8 42.1 65.4 45.1 41/— imgs/s 49M 266.43G/ —
Swin-T-GLOBAL 48.0 70.0 52.7 43.3 67.0 46.8 6/17 imgs/s 48M 2106.75G/260.48G
Swin-T-LOCAL 46.0 68.1 50.3 41.6 65.1 44.9 45/— imgs/s 48M 267.01G/ —
ConvNeXt-T 46.2 67.9 50.8 41.7 65.0 44.9 44/— imgs/s 48M 262.29G/ —
VMamba-T 48.9 70.6 53.6 43.7 67.7 46.8 35/— imgs/s 50M 271.16G/ —
IV-E Ablation Study of Different Combinations of $c_{1}$ and $c_{2}$
Table V shows the performance comparison of CSWin-T24181-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. When $c_{1}=1\times c$ and $c_{2}=0\times c$, ELFATT becomes EFFATT, and when $c_{1}=0\times c$ and $c_{2}=1\times c$, ELFATT becomes a local window-based attention mechanism which can be regarded as a simpler version of the local window-based attention mechanism used in Swin. As shown in Table V, $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ achieve better performance than other combinations of $c_{1}$ and $c_{2}$. Without using FlashAttention-2 acceleration, the difference between $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ in terms of speed is not significant. Using FlashAttention-2 acceleration, $c_{1}=0\times c$ and $c_{2}=1\times c$, and $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$ achieve close speed and are significantly faster than other combinations.
TABLE VII: The comparison of semantic segmentation performance of all methods on ADE20K. Note: The best values are in bold, ‘mAcc’ denotes mean class accuracy, and ‘mIoU’ denotes mean intersection over union. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. FLOPs are calculated using an input size of $512\times 2048$. ‘160k’ denotes the 160k fine-tuning iterations. Inference throughput is obtained using a batch size of 1 with mixed precision on a single NVIDIA H20 (96 GB) GPU. “imgs” denotes images, “nFA” denotes non-FlashAttention-2 acceleration, and “FA” denotes FlashAttention-2 acceleration.
UperNet [49] 160k
Method mAcc (%) mIoU (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-T24181-Agent 60.8 48.5 16/17 imgs/s 50M 953.60G/929.64G
CSWin-T24181-EFFATT 60.6 48.8 28/33 imgs/s 50M 1008.73G/930.34G
CSWin-T24181-ELFATT 61.2 49.6 28/32 imgs/s 50M 1014.26G/929.53G
CSWin-T24181-FLatten 61.4 49.3 25/27 imgs/s 51M 954.15G/930.19G
CSWin-T24181-GLOBAL 61.1 48.8  6/14 imgs/s 50M 2458.75G/928.67G
CSWin-T24181-LOCAL 61.1 49.6 26/28 imgs/s 50M 963.38G/928.68G
Swin-T-Agent 58.5 46.7 4/— imgs/s 61M 957.50G/ —
Swin-T-EFFATT 58.3 46.7 35/39 imgs/s 60M 981.22G/939.35G
Swin-T-ELFATT 59.3 47.7 34/38 imgs/s 62M 991.27G/943.94G
Swin-T-FLatten 57.0 44.8 35/— imgs/s 61M 944.62G/ —
Swin-T-GLOBAL 59.3 47.8 5/14 imgs/s 60M 2873.79G/937.84G
Swin-T-LOCAL 55.6 44.5 38/— imgs/s 60M 945.66G/ —
ConvNeXt-T 58.3 46.1 37/— imgs/s 60M 939.69G/ —
VMamba-T 59.3 47.9 34/— imgs/s 62M 948.78G/ —
IV-F Object Detection Performance
Table VI shows the object detection performance of all methods on MS COCO 2017. Using the Mask-RCNN [48] 1$\times$ schedule, the vanilla softmax-based attention mechanism and ELFATT significantly outperform other attention mechanisms. ELFATT exhibits slightly better performance than the vanilla softmax-based attention mechanism when using CSWin-T24181 as the backbone and significantly outperforms the vanilla softmax-based attention mechanism when using Swin-T as the backbone. Under the backbone of CSWin-T24181, ELFATT offers a 4.3x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 acceleration and is still 1.8x faster than the vanilla softmax-based attention mechanism using FlashAttention-2 acceleration. When using Swin-T as the backbone, ELFATT offers a 6.5x speedup over the vanilla softmax-based attention mechanism without using FlashAttention-2 acceleration and is still 2.6x faster than the vanilla softmax-based attention mechanism using FlashAttention-2 acceleration. Using the Mask-RCNN [48] 3$\times$ multiscale training schedule, ELFATT still achieves the best performance. The vanilla softmax-based attention mechanism still outperforms most linear attention mechanisms when using Swin-T as the backbone. In the 1$\times$ schedule, ELFATT using CSWin-T24181 as the backbone achieves close performance compared to VMamba-T and significantly outperforms ConvNeXt-T. In the 3$\times$ multiscale training schedule, ELFATT using CSWin-T24181 as the backbone outperforms ConvNeXt-T and VMamba-T in terms of object detection performance.
TABLE VIII: The comparison of inference throughput of different attention mechanisms obtained on ImageNet-1K (Platform: NVIDIA Jetson AGX Orin; Batch Size: 128; Mode: 60W (Orin MAXN)).
Method Res. FPS (FP32) Speedup ratio FPS (Mixed precision) Speedup ratio
CSWin-T24181-Agent $224^{2}$ 158 imgs/s 1.5x 295 imgs/s 1.3x
CSWin-T24181-EFFATT $224^{2}$ 156 imgs/s 1.4x 305 imgs/s 1.3x
CSWin-T24181-ELFATT $224^{2}$ 174 imgs/s 1.6x 355 imgs/s 1.6x
CSWin-T24181-FLatten $224^{2}$ 138 imgs/s 1.3x 229 imgs/s 1.0x
CSWin-T24181-GLOBAL $224^{2}$ 108 imgs/s 1.0x 298 imgs/s 1.3x
CSWin-T24181-LOCAL $224^{2}$ 167 imgs/s 1.5x 309 imgs/s 1.3x
EfficientViT-B2 $288^{2}$ 198 imgs/s 1.8x 282 imgs/s 1.2x
TABLE IX: The comparison of inference throughput of different attention mechanisms obtained on ImageNet-1K (Platform: NVIDIA Jetson Nano; Batch Size: 1; Mode: 10W (Nano MAXN)).
Method Res. FPS (FP32) Speedup ratio FPS (Mixed precision) Speedup ratio
CSWin-T24181-Agent $224^{2}$ 4.9 imgs/s 1.5x 4.5 imgs/s 1.3x
CSWin-T24181-EFFATT $224^{2}$ 5.8 imgs/s 1.8x 5.5 imgs/s 1.6x
CSWin-T24181-ELFATT $224^{2}$ 6.2 imgs/s 2.0x 5.8 imgs/s 1.7x
CSWin-T24181-FLatten $224^{2}$ 4.7 imgs/s 1.5x 4.0 imgs/s 1.1x
CSWin-T24181-GLOBAL $224^{2}$ 3.2 imgs/s 1.0x 3.5 imgs/s 1.0x
CSWin-T24181-LOCAL $224^{2}$ 5.0 imgs/s 1.6x 4.2 imgs/s 1.2x
EfficientViT-B2 $288^{2}$ 5.5 imgs/s 1.7x 5.7 imgs/s 1.6x
Figure 4: The speed comparison of different attention mechanisms across various power modes from 5W to 60W by using NVIDIA Jetson Nano or NVIDIA Jetson AGX Orin.
IV-G Semantic Segmentation Performance
Table VII shows the comparison of semantic segmentation performance of all methods on ADE20K. The vanilla softmax-based attention mechanism and ELFATT achieve close mean class accuracy (mAcc) and mean intersection over union (mIoU), and significantly outperform other attention mechanisms when using Swin-T as the backbone. ELFATT, FLatten, the vanilla softmax-based attention mechanism, and the local window-based attention mechanism [23] achieve close mean class accuracy when using CSWin-T24181 as the backbone. However, ELFATT and the local window-based attention mechanism [23] significantly outperform other attention mechanisms in terms of mean intersection over union. Without using FlashAttention-2 acceleration, ELFATT offers a nearly 5x speedup over the vanilla softmax-based attention mechanism using CSWin-T24181 as the backbone and a 7x speedup over the vanilla softmax-based attention mechanism using Swin-T as the backbone. Even using FlashAttention-2 acceleration, under the CSWin-T24181 backbone, the vanilla softmax-based attention mechanism is still 2x slower than ELFATT without using FlashAttention-2 acceleration and 2.3x slower than ELFATT using FlashAttention-2 acceleration. Under the Swin-T backbone, the vanilla softmax-based attention mechanism is still 2.4x slower than ELFATT without using FlashAttention-2 acceleration and 2.7x slower than ELFATT using FlashAttention-2 acceleration. The speed of ELFATT is almost the same as the speed of EFFATT which is a real linear attention mechanism. However, ELFATT achieves significantly higher mean class accuracy and mean intersection over union than those of EFFATT. ELFATT using CSWin-T24181 as the backbone outperforms ConvNeXt-T and VMamba-T in terms of semantic segmentation performance.
IV-H Speed Comparison on Edge GPUs
In addition to high-performance computing applications, model deployment should also be evaluated in emerging fields such as robotic vision, unmanned aerial vehicles (UAV), and autonomous driving. These scenarios typically operate under power constraints while guaranteeing maximum performance and efficiency. The NVIDIA Jetson series provides a good embedded and edge computing platform to evaluate AI model performance in such constrained environments.
The experiments were conducted to assess the model’s inference speed on NVIDIA Jetson platforms, including Jetson Nano and Jetson AGX Orin, across power modes ranging from 5W to 60W. Each model was evaluated in full-precision (FP32) and mixed-precision modes. The evaluations were performed using the ImageNet-1K dataset with a batch size of 1 for the Jetson Nano and 128 for the Jetson AGX Orin. Both devices installed the latest NVIDIA JetPack SDK. The Jetson Nano used JetPack SDK 4.6.6 and the Jetson AGX Orin utilized JetPack SDK 6.1. Each experiment was repeated 100 times and the reported results represent the average of these repetitions to ensure statistical reliability.
Tables VIII and IX show the comparison of inference speed of different attention mechanisms obtained on ImageNet-1K using an NVIDIA Jetson AGX Orin/NVIDIA Jetson Nano GPU. In both FP32 and mixed precision, under the backbone of CSWin-T24181, ELFATT achieves the highest FPS. In mixed precision, ELFATT is significantly faster than all other attention mechanisms, as shown in Tables VIII and IX. Also, from Tables VIII and IX, FLatten does not have a significant acceleration effect compared to the vanilla softmax-based attention mechanism and is even slower than the vanilla softmax-based attention mechanism in mixed precision on the edge GPUs. Because FLatten uses a lot of normalization and reshaping operations in addition to linear attention. It further shows that these kinds of linear attention are not linear for all platforms. Even compared to EfficientViT-B2 [50], to achieve similar accuracy, ELFATT is significantly faster than EfficientViT-B2 in mixed precision on the edge GPUs. Fig. 4 shows the speed comparison of different attention mechanisms across various power modes from 5W to 60W by using NVIDIA Jetson Nano or NVIDIA Jetson AGX Orin. ELFATT consistently achieves 1.6x to 2.0x speedups compared to other attention mechanisms in various power modes from 5W to 60W. Compared to EfficientViT-B2, in mixed precision, ELFATT performs on par with EfficientViT-B2 in low power conditions, such as the 5W mode. As power increases, the speed growth rate of ELFATT surpasses that of EfficientViT-B2, especially in high-power modes like 50W and 60W, where it demonstrates a significant performance advantage. In mixed precision, ELFATT outperforms all other models in terms of speed across all tested power modes, establishing itself as the fastest option in both high-performance and power constrained scenarios.
V Conclusion
In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism for ViTs to achieve low memory I/O operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT is compatible with FlashAttention. Using FlashAttention-2 acceleration, ELFATT still offers 2-3x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT without using FlashAttention-2 acceleration is even faster than the vanilla softmax-based attention mechanism using FlashAttention-2 acceleration on both low-resolution and high-resolution vision datasets, which shows great potential of ELFATT. What’s more, on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W.