### v4

# ELFATT: Efficient Linear Fast Attention for Vision Transformers

Abstract.
The attention mechanism is the key to the success of transformers in different machine learning tasks. However, the quadratic complexity with respect to the sequence length of the vanilla softmax-based attention mechanism becomes the major bottleneck for the application of long sequence tasks, such as vision tasks. Although various efficient linear attention mechanisms have been proposed, they need to sacrifice performance to achieve high efficiency. What’s more, memory-efficient methods, such as FlashAttention-1-3, still have quadratic computation complexity which can be further improved. In this paper, we propose a novel efficient linear fast attention (ELFATT) mechanism to achieve low memory input/output operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over the vanilla softmax-based attention mechanism in high-resolution vision tasks without losing performance. ELFATT is FlashAttention friendly. Using FlashAttention-2 acceleration, ELFATT still offers 2-3x speedups over the vanilla softmax-based attention mechanism on high-resolution vision tasks without losing performance. Even in some non-vision tasks of long-range arena, ELFATT still achieves leading performance and offers 1.2-2.3x speedups over FlashAttention-2. Even on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W. Furthermore, ELFATT can be used to enhance and accelerate diffusion tasks directly without training.
Efficient Attention Mechanism; Linear Approximation; Vision Transformer
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3754825isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Machine learningccs: Theory of computation Design and analysis of algorithmsccs: Computing methodologies Neural networksccs: Computing methodologies Computer vision
1. Introduction
Transformers have achieved great success in large language models (ChatGPT (Achiam et al., 2023) and Llama (Grattafiori et al., 2024)) and large vision models (SAM (Kirillov et al., 2023) and Sora (Brooks et al., 2024)). The core technique of the transformer, the vanilla softmax-based attention (VaniATT) mechanism, is capable of capturing the relationship between any two tokens (Wu et al., 2024). In complex tasks, the sequence length becomes longer and longer, usually much longer than the embedding dimension. The softmax operation after the multiplication of query and key matrices makes VaniATT quadratic computational complexity with respect to the sequence length. It has become a main obstacle to computational efficiency.
Acceleration methods to speed up the attention computation can be classified into two categories: (1) memory-efficient methods; (2) computation-efficient methods. Memory-efficient methods focus on optimizing memory input/output (I/O) operations to achieve almost linear complexity (Dao et al., 2022; Dao, 2024; Shah et al., 2024; Ramapuram et al., 2025). FlashAttention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) and FlashSigmoid (Ramapuram et al., 2025) are representatives of memory-efficient methods. Computation-efficient methods focus on minimizing the computation bound of the attention computation by using linear approximations based on different kernel methods (Han et al., 2023; Choromanski et al., 2021; Wortsman et al., 2023; Katharopoulos et al., 2020; Bolya et al., 2023; Lu et al., 2021; Peng et al., 2021; Qin et al., 2022; Shen et al., 2021; Han et al., 2025; Ran et al., 2025) or low-rank decomposition (Han et al., 2022; Xiong et al., 2021; Lu et al., 2021; Wu et al., 2024), and sparse computation (Liu et al., 2021; Dong et al., 2022; Wang et al., 2020a; Zhao et al., 2019; Beltagy et al., 2020; Child et al., 2019; Tay et al., 2020; Zaheer et al., 2020; Kitaev et al., 2020). However, memory-efficient methods still have quadratic computational complexity. And computation-efficient methods usually have lower performance compared to VaniATT (Liu et al., 2023).
In this paper, we propose a novel efficient linear fast attention (ELFATT [^1] The ELFATT: Efficient Linear Fast Attention for Vision Transformers methodology disclosed in this work is protected under intellectual property rights, including the following patent applications and granted patents: U.S. Non-Provisional Patent Application (No. 19/011,779 filed on 7 January 2025), Hong Kong Short-Term Patent (No. HK30118021A2 granted on 30 May 2025), and Chinese Invention Patent Application (No. 202510108088.7 filed on 23 January 2025). Commercial inquiries should be directed to kt@innocimda.com.) mechanism that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to VaniATT. The core idea of ELFATT is the combination of sparse computation with a linear approximation. Each ELFATT block has two parallel attention heads. One head is used to compute sparse blockify attention to introduce inductive biases, and the other one head is used to compute global linear attention to capture long-range dependencies. Both heads have almost linear complexity, and the sparse blockify attention head can be further speeded up by using the FlashAttention (Dao et al., 2022; Dao, 2024; Shah et al., 2024) mechanisms to reduce memory I/O operations. ELFATT is evaluated on different vision tasks: image classification, semantic segmentation, object detection, and diffusion. Furthermore, ELFATT is also evaluated on edge GPUs and in the long-range arena. Compared to state-of-the-art memory-efficient acceleration methods and computation-efficient acceleration methods, ELFATT inherits advantages of both two kinds of methods: noninferior performance compared to VaniATT, low memory I/O operations, and linear computational complexity. In summary, this paper has the following contributions.
  • (i)
    A novel efficient linear fast attention (ELFATT) mechanism is proposed that has low memory I/O operations and linear computational complexity and maintains noninferior performance compared to VaniATT.
  • (ii)
    ELFATT is the first sparse blockify attention + global linear attention design for attention acceleration which is significantly different from the existing hierarchical sparse attention mechanisms.
  • (iii)
    The relationship between ELFATT and VaniATT is analyzed and given. An upper bound is given for the use of ELFATT to approximate VaniATT. To the best of our knowledge, we are the first to provide a theoretical bound analysis to sparse attention.
  • (iv)
    ELFATT offers 4-7x speedups over VaniATT without using FlashAttention-2 and 2-3x speedups over VaniATT using FlashAttention-2 on high-resolution vision tasks.
  • (v)
    ELFATT can be used to accelerate non-vision tasks with 1K-4K sequence lengths and offers 1.2-2.3x speedups over FlashAttention-2.
  • (vi)
    ELFATT offers 1.6x to 2.0x speedups over state-of-the-art attention mechanisms in various power modes from 5W to 60W on edge GPUs.
  • (vii)
    ELFATT can be used to enhance and accelerate diffusion tasks directly without training.
2. Related Work
Since VaniATT has quadratic complexity in both time and memory, depending on whether the focus is on optimizing memory complexity or time complexity, attention acceleration methods can be categorized into the following two types: (A) memory-efficient attention acceleration methods; (B) computation-efficient attention acceleration methods.
2.1. Memory-Efficient Attention Acceleration Methods
Memory-efficient attention acceleration methods focus on optimizing quadratic memory complexity to almost linear complexity. One of the most representative methods is FlashAttention (Dao et al., 2022), which minimizes memory I/O access for the softmax computation of the product of query and key matrices to improve the utilization rate of GPUs to achieve almost linear memory complexity (Dao et al., 2022; Wu et al., 2024). FlashAttention-2 further reduces the number of floating point operations (FLOPs) of non-matrix multiplication, parallelizes both forward and backward processes according to the sequence length, and reduces the I/O access of shared memory (Dao, 2024). FlashAttention-3 is proposed to fully utilize the performance of new Hopper GPUs by introducing warp-specialization, interleave block-wise matrix multiplication and softmax operations, and the support of FP8 precision (Shah et al., 2024). FlashSigmoid investigates the feasibility of using the sigmoid function to replace the softmax function in attention computation, proposes a new regularity method for sigmoid attention to stabilize training, and introduces a memory-efficient version based on FlashAttention-2 (Ramapuram et al., 2025). However, these memory-efficient attention acceleration methods are usually hard aware and can not support all kinds of GPUs and their computation complexity is still quadratic.
2.2. Computation-Efficient Attention Acceleration Methods
Different to hard aware memory-efficient attention acceleration methods, computation-efficient attention acceleration methods focus on optimizing computational complexity by linear approximations and usually can be categorized into two classes: (a) kernel methods; (b) sparse methods.
Kernel methods usually change the order of nonlinear normalization and multiplication of the query matrix and key matrix to achieve linear computational complexity. FLatten (Han et al., 2023) introduces a cubic rectified linear unit (ReLU) (Nair and Hinton, 2010) based feature map and uses it to perform nonlinear normalization on the query matrix and key matrix, respectively, before the attention computation. The linear transformer (Katharopoulos et al., 2020) introduces a feature map based on the exponential linear unit (ELU) (Clevert, 2015) activation function. cosFormer (Qin et al., 2022) introduces a cos-based nonlinear feature map to obtain a linear approximation of VaniATT. Efficient attention (Shen et al., 2021) performs softmax normalization on the query matrix and the key matrix, respectively, to achieve linear complexity. However, this linearization introduces noise from two aspects: (1) Separated softmax normalization will reduce the similarity of elements in the query matrix and key matrix that are both negative at the same position; (2) Large negative values are suppressed to small positive values. The above two types of noise result in decreased discrimination between tokens and incorrect concentration of attention. Agent (Han et al., 2025) introduces a small agent matrix, which is obtained by performing the pooling operation on the query matrix. This agent matrix is used as the auxiliary key matrix to be multiplied with the query matrix to reduce its dimenison before the softmax normalization. Similarly, it is also used to reduce the dimension of the key matrix before its softmax normalization. This agent matrix serves as a bridge between two independent softmax normalization processes to reduce noise. Another type of kernel methods introduce the low-rank decomposition (Xu et al., 2024b, a; Xu and Wei, 2024). They usually take the softmax normalization of the product of query and key matrices as a whole for decomposition to find appropriate feature maps to reduce complexity (Wu et al., 2024). Nyströmformer (Xiong et al., 2021) introduces Nyström approximation to perform low-rank decomposition of the softmax normalization of the product of query and key matrices. For calculating an attention score matrix, it needs to perform an iterative inverse approximation and operate softmax normalization three times (each softmax normalization still involves a product of two matrices), which makes its acceleration effect not significant. Skeleton decomposition-based self-attention uses a simplified inverse approximation method based on the permuted diagonal matrix (Han et al., 2022). Interactive multihead self-attention (iMHSA) introduces several linear layers to perform head interactions instead of performing an iterative inverse approximation (Kang et al., 2024). CUR decomposition based self-attention (CURSA) (Wu et al., 2024) introduces the CUR decomposition to replace the Nyström approximation and reduces the number of matrix multiplication. Mamba (Gu and Dao, 2024) and its variants (Dao and Gu, 2024; Liu et al., 2024; Zhu et al., 2024) are also a kind of kernelized attention mechanism.
Sparse methods usually separate the original sequences into smaller blocks using sliding windows and perform attention computation within these blocks to reduce complexity. Each token in a sequence is only affected by several tokens, not all tokens; hence, this kind of sparse method is also called the local attention mechanism. Swin (Liu et al., 2021) and Longformer (Beltagy et al., 2020) are pioneers in introducing sliding windows into attention computation for vision tasks and language tasks, respectively. To address the information loss introduced by the local windows, Swin introduces a shifted window mechanism to capture cross-block information, while Longformer selects some tokens as global tokens, which have effects on all tokens, to compensate for the global information loss. NesT (Zhang et al., 2022) further simplifies the shifted window mechanism through simple spatial operations. CSWin (Dong et al., 2022) introduces a cross-shaped window mechanism for 2D sequences. The cross-shaped window mechanism separates a 2D sequence into horizontal and vertical stripes, respectively, and performs parallel attention computation within these stripes in horizontal and vertical directions, which can simultaneously obtain local inductive biases and cross-block (global) information. In addition to using the deterministic global token selection mechanism of Longformer, Big Bird (Zaheer et al., 2020) also randomly selects some global tokens to enhance performance. Similarly, the sparse transformer (Child et al., 2019) proposed by OpenAI also introduces several fixed-step tokens, which is similar to dilated convolution (Yu, 2015) in convolutional neural networks (CNNs) to capture long-range information. Reformer (Kitaev et al., 2020) finds the nearest neighbors for each token to calculate local attention scores using the locality-sensitive-hashing (LSH) algorithm.
3. Methods
3.1. Vanilla Softmax-Based Attention Mechanism
For clarity, we assume that all vectors appear in this paper are row vectors. Scaling and normalization are omitted in this paper for convenience. For any two same sized tokens $\textit{{x}}\in\mathbb{R}^{c}$ and $\textit{{y}}\in\mathbb{R}^{c}$, their attention similarity score can be calculated as follows,
(1) $a={\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right),$
where ${\rm exp}\left(\cdot\right)$ is an element-wise exponential function and is conducted after the inner product of two tokens. For given two $m$-tuples $\{\textit{{x}}_{1},\dots,\textit{{x}}_{m}\}$ and $\{\textit{{y}}_{1},\dots,\textit{{y}}_{m}\}$ with $\textit{{x}}_{i},\textit{{y}}_{i}\in\mathbb{R}^{c}$ and $i=1,2,\dots,m$, we need to calculate $m^{2}$ pairs of attention similarity scores. Currently, a long sequence length $m$ is preferred in large models and $m\gg c$, therefore, the computational complexity of VaniATT is quadratic with respect to the sequence length $m$.
3.2. General Attention Mechanism
Eq. (1) can be rewritten as a more general form (Katharopoulos et al., 2020) as follows,
(2) $a={\rm sim}\left(\textit{{x}},\textit{{y}}\right),$
where ${\rm sim}(\cdot)$ is a non-negative similarity function and it satisfies the definition of a kernel function $G(\textit{{x}},\textit{{y}}):\mathbb{R}^{c}\times\mathbb{R}^{c}\rightarrow\mathbb{R}_{+}$ with +=$\mathbb{R}_{+}=\{x\in\mathbb{R}:x\geq 0\}$.
3.3. Kernelized Attention Mechanism
If we have such a kernel with a non-negative feature map $\phi$, Eq. (2) can be written as follows,
(3) $a=\phi\left(\textit{{x}}\right)\phi\left(\textit{{y}}\right)^{\top}.$
Performer (Choromanski et al., 2021) has proved that one of the best choices of the non-negative feature map $\phi$ for Eq. (3) is $\rm exp\left(\cdot\right)$. Performer obtains an exact alternative of Eq. (1) using random feature maps as follows,
(4) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)=\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}\right)\right],$
where $e\left(\textit{{x}}\right)={\rm exp}\left(-\frac{\|\textit{{x}}\|^{2}}{2}\right)$, $e\left(\textit{{y}}\right)={\rm exp}\left(-\frac{\|\textit{{y}}\|^{2}}{2}\right)$, $\textbf{0}_{c}\in\mathbb{R}^{c}$ is the zero vector, and $\textbf{I}_{c}\in\mathbb{R}^{c\times c}$ is the identity matrix. If the attention scores for all pairs of x and $\textit{{y}}_{i}$ ($i=1,2,\dots,m$) using Eq. (4) are obtained, after the following normalization $e\left(\textit{{x}}\right)$ will be canceled.
$\begin{split}&\frac{{\rm exp}\left(\textit{{x}}\textit{{y}}_{i}^{\top}\right)}{\sum_{j=1}^{m}{{\rm exp}\left(\textit{{x}}\textit{{y}}_{j}^{\top}\right)}}\\ &=\frac{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{i}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}_{i}\right)\right]}{\sum_{j=1}^{m}{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{j}^{\top}\right)e\left(\textit{{x}}\right)e\left(\textit{{y}}_{j}\right)\right]}}\\ &=\frac{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{i}^{\top}\right)e\left(\textit{{y}}_{i}\right)\right]}{\sum_{j=1}^{m}{\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}_{j}^{\top}\right)e\left(\textit{{y}}_{j}\right)\right]}}.\end{split}$
Hence, $e\left(\textit{{x}}\right)$ has no effect on the final attention score. Eq. (4) can be written as follows,
(5) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)e\left(\textit{{y}}\right)\right].$
After the normalization, the left and right-hand sides of Eq. (5) will be equal. Eq. (4) only holds when taking the sum of an infinite number of random vectors $\omega$. To avoid performing summation of infinite terms, Performer samples $c\times\log(c)$ random vectors $\omega$ to ensure a low approximation error. If the change of $e\left(\textit{{y}}\right)$ is much smaller than ${\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)$, Eq. (5) can be further simplified and approximated as follows,
(6) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\mathbb{E}_{\omega\sim\mathcal{N}(\textbf{0}_{c},\textbf{I}_{c})}\left[{\rm exp}\left(\omega\textit{{x}}^{\top}\right){\rm exp}\left(\omega\textit{{y}}^{\top}\right)\right].$
Efficient attention (EFFATT) (Shen et al., 2021) is a special case of Eq. (6). It simplifies $\omega$ to a one-hot vector and only uses $c$ one-hot vectors in Eq. (6) to obtain an approximation as follows,
(7) ${\rm exp}\left(\textit{{x}}\textit{{y}}^{\top}\right)\approx\frac{1}{c}{\rm exp}\left(\textit{{x}}\right){\rm exp}\left(\textit{{y}}\right)^{\top}.$
Its approximation error was studied and obtained in (Wu et al., 2024). Eq. (7) has a problem of concentration reduction of attention maps (Wu et al., 2024) and Performer also has this problem when the number of $\omega$ sampled is too small. When the number of $\omega$ sampled is too large, Performer may be slower than VaniATT.
3.4. Efficient Linear Fast Attention Mechanism
To address the defects of Eq. (7), in this paper, we propose a novel attention mechanism as follows,
(8) $\begin{split}{\rm exp}&\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$
where $\textit{{Q}}=[\bar{\textit{{Q}}},\tilde{\textit{{Q}}}]\in\mathbb{R}^{m\times c}$, $\textit{{K}}=[\bar{\textit{{K}}},\tilde{\textit{{K}}}]\in\mathbb{R}^{m\times c}$, $\textit{{V}}=[\bar{\textit{{V}}},\tilde{\textit{{V}}}]\in\mathbb{R}^{m\times c}$, $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{Q}}}\in\mathbb{R}^{m\times c_{2}}$, $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{K}}}\in\mathbb{R}^{m\times c_{2}}$, $\bar{\textit{{V}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{V}}}\in\mathbb{R}^{m\times c_{2}}$, $c=c_{1}+c_{2}$, $f\left(\cdot\right)$ is a blockify function to separate a matrix with a size of $m\times c_{2}$ into $b$ blocks (each block has a size of $(m/b)\times c_{2}$), and $g\left(\cdot\right)$ is an unblockify function to unblock $b$ blocks to a single matrix with a size of $m\times c_{2}$.
Eq. (8) denotes the single-head VaniATT mechanism approximated by using ELFATT. From the right-hand side of Eq. (8), each ELFATT attention process consists of two heads. Hence, ELFATT can directly become an approximation of the double-head VaniATT mechanism as follows,
(9) $\begin{split}\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\approx&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right].\end{split}$
Since each ELFATT block corresponds to two parallel heads as Fig. 1 shows, $s$ ($s>0$) ELFATT will be needed for the approximation of $2\times s$ heads of VaniATT according to Eq. (9). For the approximation of $2\times s-1$ heads of VaniATT, $2\times s-1$ ELFATT will be needed according to Eq. (8). The approximation error bound analysis of Eqs. (8) and (9) is available in Section A and Fig. 2 of Appendix.
3.5. Positional Encoding
Different positional encoding mechanisms such as absolute positional encoding (Vaswani et al., 2017), relative positional encoding (Shaw et al., 2018), and conditional positional encoding (Chu et al., 2023), have been proposed to make use of the ordering information of sequences in vision transformers (Liu et al., 2021; Wu et al., 2021; Chu et al., 2023). Among these positional encoding mechanisms, the locally enhanced positional encoding (LePE) (Dong et al., 2022) mechanism shows more powerful local positional information enhancement and brings a higher performance gain for vision transformers. Hence, LePE is selected as the positional encoding mechanism for ELFATT. After introducing LePE, Eq. (8) will become as follows,
(10) $\begin{split}{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}+&L\left(\textit{{V}}\right)\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)\right],\end{split}$
(a) VaniATT
(b) ELFATT
Figure 1. The comparison of VaniATT and ELFATT. For an input embedding matrix $\textit{{H}}\in\mathbb{R}^{m\times n}$, each ELFATT block will process it in two parallel attention heads to obtain two new embedding matrices, $\bar{\textit{{H}}}\in\mathbb{R}^{m\times c_{1}}$ and $\tilde{\textit{{H}}}\in\mathbb{R}^{m\times c_{2}}$, respectively. After a concatenation operation ${\rm cat}$, the ELFATT block will output the updated embedding matrix $\textit{{H}}=[\bar{\textit{{H}}},\tilde{\textit{{H}}}]\in\mathbb{R}^{m\times c}$.
Eq. (9) will become as follows,
(11) $\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}+L\left(\tilde{\textit{{V}}}\right)\right]\approx\\ &\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)\right],\end{split}$
where $L(\cdot)$ denotes a depthwise convolution operation with a kernel size of 3. Fig. 1 shows the comparison of VaniATT and ELFATT. $\mathcal{W}_{\textit{{Q}}}$, $\mathcal{W}_{\textit{{K}}}$, and $\mathcal{W}_{\textit{{V}}}\in\mathbb{R}^{n\times c}$ are three projection matrices. For an input embedding matrix $\textit{{H}}\in\mathbb{R}^{m\times n}$, each ELFATT block will process it in two parallel attention heads to obtain two new embedding matrices, $\bar{\textit{{H}}}\in\mathbb{R}^{m\times c_{1}}$ and $\tilde{\textit{{H}}}\in\mathbb{R}^{m\times c_{2}}$, respectively. After a concatenation operation, the ELFATT block will output the updated embedding matrix $\textit{{H}}=[\bar{\textit{{H}}},\tilde{\textit{{H}}}]\in\mathbb{R}^{m\times c}$.
Table 1. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of different methods on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 for tiny models, and 256/32 for base models with a resolution of $224^{2}$/$384^{2}$ using mixed precision on a single NVIDIA H20 (96 GB) GPU. “—” denotes the corresponding method cannot be accelerated by FlashAttention-2. “Res.” denotes resolution, “imgs” denotes images, and “nFA/FA” denotes without/with using FlashAttention-2.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-B
Agent $224^{2}$ 84.7 930/994 imgs/s 73M 14.92G/14.49G
EFFATT $224^{2}$ 84.4 985/1059 imgs/s 73M 14.98G/14.53G
ELFATT $224^{2}$ 84.7 1000/1187 imgs/s 73M 15.47G/14.46G
FLatten $224^{2}$ 84.5 814/864 imgs/s 75M 14.96G/14.52G
GLOBAL $224^{2}$ 84.7 478/879 imgs/s 73M 22.33G/14.39G
LOCAL $224^{2}$ 84.4 941/1037 imgs/s 73M 15.03G/14.39G
Agent $384^{2}$ 85.8 246/276 imgs/s 73M 46.33G/42.57G
EFFATT $384^{2}$ 85.7 294/331 imgs/s 73M 46.57G/42.69G
ELFATT $384^{2}$ 85.8 272/355 imgs/s 73M 51.21G/42.48G
FLatten $384^{2}$ 85.5 238/266 imgs/s 78M 46.43G/42.67G
GLOBAL $384^{2}$ 85.9 82/201 imgs/s 73M 110.89G/42.28G
LOCAL $384^{2}$ 85.5 276/323 imgs/s 73M 47.06G/42.28G
CSWin-T
Agent $224^{2}$ 83.1 2297/2425 imgs/s 20M 4.31G/4.14G
EFFATT $224^{2}$ 82.6 2394/2526 imgs/s 20M 4.35G/4.17G
ELFATT $224^{2}$ 83.1 2603/2856 imgs/s 20M 4.44G/4.13G
FLatten $224^{2}$ 83.1 1934/2025 imgs/s 21M 4.34G/4.16G
GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
LOCAL $224^{2}$ 82.7 2330/2519 imgs/s 20M 4.36G/4.09G
Swin-B
Agent $224^{2}$ 84.0 1367/ —   imgs/s 88M 15.44G/   —
EFFATT $224^{2}$ 84.2 1439/1536 imgs/s 88M 15.69G/15.33G
ELFATT $224^{2}$ 84.5 1314/1497 imgs/s 91M 16.46G/15.68G
FLatten $224^{2}$ 83.8 1226/ —   imgs/s 89M 15.46G/   —
GLOBAL $224^{2}$ 84.2 743/1325 imgs/s 88M 21.57G/15.19G
LOCAL $224^{2}$ 83.5 1351/ —   imgs/s 88M 15.47G/   —
Agent $384^{2}$ 84.9 372/ — imgs/s 88M 46.34G/   —
EFFATT $384^{2}$ 85.3 433/481 imgs/s 88M 48.18G/45.04G
ELFATT $384^{2}$ 85.5 372/457 imgs/s 91M 52.79G/46.08G
FLatten $384^{2}$ 85.0 353/ — imgs/s 91M 46.49G/   —
GLOBAL $384^{2}$ 85.3 134/313 imgs/s 88M 99.76G/44.64G
LOCAL $384^{2}$ 84.5 357/ — imgs/s 88M 47.19G/   —
Swin-T
Agent $224^{2}$ 82.6 2847/ —   imgs/s 29M 4.53G/ —
EFFATT $224^{2}$ 82.1 3165/3282 imgs/s 28M 4.55G/4.45G
ELFATT $224^{2}$ 82.7 2884/3159 imgs/s 30M 4.99G/4.67G
FLatten $224^{2}$ 82.1 2502/ —   imgs/s 29M 4.50G/ —
GLOBAL $224^{2}$ 82.4 1269/2571 imgs/s 28M 8.81G/4.38G
LOCAL $224^{2}$ 81.4 2943/ —   imgs/s 28M 4.51G/ —
Others
ConvNeXt-T $224^{2}$ 82.1 3911/ —   imgs/s 29M 4.47G/ —
VMamba-T $224^{2}$ 82.5 1837/ —   imgs/s 30M 4.84G/ —
4. Complexity Analysis
According to Eqs. (10) and (11), the complexity of ELFATT can be divided into two parts: (a) the global linear attention head ${\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right)$; (b) the sparse blockify attention head $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)$. The complexity of the global linear attention head is $O(m\times c_{1}^{2})$ in the case of $m>c_{1}$, and $O(m^{2}\times c_{1})$ in the case of $m\leq c_{1}$. The complexity of the sparse blockify attention head is $O((m^{2}/b)\times c_{2})$. If $m/b\ll m$, the sparse blockify attention head will also achieve almost linear complexity.
5. Experiments and Results
We evaluated ELFATT in commonly used vision tasks: image classification (ImageNet-1K (Russakovsky et al., 2015)), semantic segmentation (ADE20K (Zhou et al., 2017)), and object detection (MS COCO 2017 (Lin et al., 2015)). We compared ELFATT with VaniATT (Vaswani et al., 2017), the memory-efficient attention mechanism (FlashAttention-2 (Dao, 2024)), local window-based attention mechanisms (Swin (Liu et al., 2021) and CSWin (Dong et al., 2022)), and kernel-based attention mechanisms (Agent (Han et al., 2025), EFFATT (Shen et al., 2021), and FLatten (Han et al., 2023)). The backbone ViT architectures used to evaluate the different attention mechanisms are: Swin-T (Liu et al., 2021), Swin-B (Liu et al., 2021), CSWin-T24181 (CSWin-T) (Han et al., 2023), and CSWin-B36292 (CSWin-B) (Han et al., 2023). The original Swin-T, Swin-B, CSWin-T, and CSWin-B are called Swin-T-LOCAL, Swin-B-LOCAL, CSWin-T-LOCAL, and CSWin-B-LOCAL, respectively. The backbones after replacing local window-based attention mechanisms with VaniATT (Vaswani et al., 2017) are called Swin-T-GLOBAL, Swin-B-GLOBAL, CSWin-T-GLOBAL, and CSWin-B-GLOBAL, respectively. The backbones after replacing local window-based attention mechanisms with kernel-based attention mechanisms (Agent (Han et al., 2025), EFFATT (Shen et al., 2021), and FLatten (Han et al., 2023)) are called Swin-T-Agent, Swin-T-EFFATT, Swin-T-FLatten, Swin-B-Agent, Swin-B-EFFATT, Swin-B-FLatten, CSWin-T-Agent, CSWin-T-EFFATT, CSWin-T-FLatten, CSWin-B-Agent, CSWin-B-EFFATT, and CSWin-B-FLatten, respectively. The backbones after replacing local window-based attention mechanisms with ELFATT are called Swin-T-ELFATT, Swin-B-ELFATT, CSWin-T-ELFATT, and CSWin-B-ELFATT, respectively. For the ImageNet-1K image classification task, we used the same training settings and data augmentation methods from (Dong et al., 2022) to train all models from scratch using mixed precision. For the ADE20K semantic segmentation task (Zhou et al., 2017) and the MS COCO 2017 object detection task (Lin et al., 2015), we used the same training settings and data augmentation methods from (Liu et al., 2024) to fine-tune the weights of all models obtained from the ImageNet-1K image classification task using mixed precision. We also compared ELFATT with ConvNeXt-T (Liu et al., 2022a) and VMamba-T (Liu et al., 2024) on ImageNet-1K, ADE20K, and MS COCO 2017. The experiments of ImageNet-1K, ADE20K, and MS COCO 2017 were carried out on 8 NVIDIA vGPU (32 GB) GPUs. Inference throughput comparison experiments were carried out on 1 NVIDIA H20 (96 GB) GPU. Besides high-performance computing applications, we also compared ELFATT with state-of-the-art attention mechanisms on edge GPUs (NVIDIA Jetson AGX Orin/NVIDIA Jetson Nano). Furthermore, we investigated how ELFATT can be used to accelerate diffusion tasks. To validate the performance of ELFATT on some non-vision long sequence tasks, we introduced long-range arena (LRA) benchmark (Tay et al., 2021) and compared ELFATT with VaniATT (Vaswani et al., 2017) and efficient attention mechanisms: Linformer (Wang et al., 2020a), Nyströmformer (Xiong et al., 2021), Primal (Chen et al., 2023b), and Reformer (Kitaev et al., 2020). We followed the training protocol and the settings of (Chen et al., 2023b) and carried out the experiments on 1 NVIDIA Tesla A100 (40GB) GPU. The details of ablation studies, experiment settings of edge GPUs, and experiment settings and ablation studies of diffusion acceleration are available in Sections B, C, D, E, F, G, and H of Appendix. We used FlashAttention-2 (Dao, 2024) to speed up all models that are compatible with FlashAttention. The PyTorch implementation of ELFATT, including the detailed architecture specifications of Swin-T-ELFATT, Swin-B-ELFATT, CSWin-T-ELFATT and CSWin-B-ELFATT, is available at this URL.
5.1. Image Classification Performance
Table 1 shows the performance comparison of different attention mechanisms on ImageNet-1K. From Table 1, VaniATT (Swin-T-GLOBAL and CSWin-T-GLOBAL) can outperform most linear attention mechanisms using the same architecture. Only Agent and ELFATT achieve the most close performance compared to VaniATT in this paper. As to the inference speed comparison, the proposed ELFATT achieves the highest inference throughput (frame per second, FPS) than all other attention mechanisms when using CSWin-T as the backbone. ELFATT achieves almost the same speed as EFFATT and is significantly faster than other attention mechanisms when using Swin-T as the backbone. ELFATT offers almost a 2x speedup over VaniATT without using FlashAttention-2. VaniATT using FlashAttention-2 is still 0.1-0.2x slower than ELFATT without using FlashAttention-2. With the use of FlashAttention-2 to optimize memory operations, ELFATT can be further accelerated and offers 1.2-1.3x speedups over VaniATT. Fig. 3 in Appendix shows the visual comparison of different attention mechanisms using CSWin-T as the backbone. In Fig. 3 (a), only Agent, ELFATT, and VaniATT (GLOBAL) focus mainly on the true object: American Staffordshire Terrier. In Fig. 3 (b), only ELFATT and GLOBAL focus mainly on the true object: Hartebeest. In Fig. 3 (c), only EFFATT, ELFATT, GLOBAL, and LOCAL focus mainly on the true object: Zebra. In Fig. 3 (d), only ELFATT and GLOBAL focus mainly on the true object: Library. In Fig. 3 (e), all methods focus accurately on the true object: Snail. In Fig. 3 (f), all methods focus accurately on the true object: Black-Footed Ferret, ELFATT covers more areas of the true object. In Fig. 3 (g), all methods also focus accurately on the true object: Scottish Deerhound, ELFATT also covers more areas of the true object. In Fig. 3 (h), only ELFATT focuses mainly on the true object: Pineapple. To sum up, ELFATT shows a more focused attention map on the ground-truth object. For the larger backbone, CSWin-B, ELFATT still achieves state-of-the-art (SOTA) performance. ELFATT offers a 2.1/1.4x speedup over VaniATT with/without using FlashAttention-2 for an input resolution of $224^{2}$ and a 3.3/1.8x speedup over VaniATT with/without using FlashAttention-2 for an input resolution of $384^{2}$. In addition, ELFATT achieves significantly leading performance under the backbone of Swin-B, offering a 1.8/2.8x speedup over VaniATT without using FlashAttention-2 for an input resolution of $224^{2}$/$384^{2}$. After using FlashAttention-2, it matches the speed of EFFATT which is a real linear attention mechanism. Also, as shown in Table 1, it can be seen that although some methods have lower FLOPs than ELFATT, their actual speed is slower than ELFATT which is consistent with the observations and conclusions of (Chen et al., 2023a). Fig. 4 in Appendix shows accuracy-efficiency curves obtained by different attention mechanisms on ImageNet-1K using different bacbones. Compared to VaniATT, ELFATT offers 2.2-4.3x speedups for the backbone of CSWin and 2.5-3.4x speedups for the backbone of Swin to achieve similar accuracy. Compared to FlashAttention-2, ELFATT offers 1.3-1.8x speedups for the backbone of CSWin and 1.2-1.5x speedups for the backbone of Swin to achieve similar accuracy. ELFATT is significantly faster than all comparison methods in achieving similar accuracy using different backbones.
Table 2. The comparison of semantic segmentation performance of all methods on ADE20K. Note: ‘mAcc’ denotes mean class accuracy and ‘mIoU’ denotes mean intersection over union. FLOPs are calculated using an input size of 512×2048512$\times$ 2048512 × 2048. ‘160k’ denotes the 160k fine-tuning iterations. Inference throughput is obtained using a batch size of 1 with mixed precision on a single NVIDIA H20 (96 GB) GPU.
UperNet (Xiao et al., 2018) 160k
Method mAcc mIoU FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-T
Agent 60.8 48.5 16/17 imgs/s 50M 953.60G/929.64G
EFFATT 60.6 48.8 28/33 imgs/s 50M 1008.73G/930.34G
ELFATT 61.2 49.6 28/32 imgs/s 50M 1014.26G/929.53G
FLatten 61.4 49.3 25/27 imgs/s 51M 954.15G/930.19G
GLOBAL 61.1 48.8  6/14 imgs/s 50M 2458.75G/928.67G
LOCAL 61.1 49.6 26/28 imgs/s 50M 963.38G/928.68G
Swin-T
Agent 58.5 46.7 4/— imgs/s 61M 957.50G/ —
EFFATT 58.3 46.7 35/39 imgs/s 60M 981.22G/939.35G
ELFATT 59.3 47.7 34/38 imgs/s 62M 991.27G/943.94G
FLatten 57.0 44.8 35/— imgs/s 61M 944.62G/ —
GLOBAL 59.3 47.8 5/14 imgs/s 60M 2873.79G/937.84G
LOCAL 55.6 44.5 38/— imgs/s 60M 945.66G/ —
Others
ConvNeXt-T 58.3 46.1 37/— imgs/s 60M 939.69G/ —
VMamba-T 59.3 47.9 34/— imgs/s 62M 948.78G/ —
Table 3. The comparison of test accuracy of different attention mechanisms on LRA.
Dataset (Seqlen) ELFATT Linformer Nyström Primal Reformer VaniATT
ListOps (2K) 38.1 37.3 37.2 37.3 19.1 37.1
Text (4K) 64.8 55.9 65.5 61.2 64.9 65.0
Retrieval (4K) 81.3 79.4 79.6 77.8 78.6 79.4
Image (1K) 45.6 37.8 41.6 43.0 43.3 38.2
Pathfinder (1K) 72.3 67.6 70.9 68.3 69.4 74.2
Avg Acc. (%) 60.4 55.6 59.0 57.5 55.1 58.8
5.2. Semantic Segmentation Performance
Table 2 shows the comparison of semantic segmentation performance of all methods on ADE20K. VaniATT and ELFATT achieve close mean class accuracy (mAcc) and mean intersection over union (mIoU), and significantly outperform other attention mechanisms when using Swin-T as the backbone. ELFATT and FLatten achieve close mAcc when using CSWin-T as the backbone. However, ELFATT and the local window-based attention mechanism (Dong et al., 2022) significantly outperform other attention mechanisms in terms of mIoU. Without using FlashAttention-2, ELFATT offers a nearly 5x speedup over VaniATT using CSWin-T as the backbone and a 7x speedup over VaniATT using Swin-T as the backbone. Even using FlashAttention-2, under the CSWin-T backbone, VaniATT is still 2x slower than ELFATT without using FlashAttention-2 and 2.3x slower than ELFATT using FlashAttention-2. Under the Swin-T backbone, VaniATT is still 2.4x slower than ELFATT without using FlashAttention-2 and 2.7x slower than ELFATT using FlashAttention-2. The speed of ELFATT is almost the same as the speed of EFFATT which is a real linear attention mechanism. However, ELFATT achieves significantly higher mAcc and mIoU than those of EFFATT. ELFATT outperforms ConvNeXt-T and VMamba-T in terms of segmentation performance.
Table 4. The comparison of running time (s) per 1K training steps and peak memory consumption of different attention mechanisms on LRA on a single NVIDIA H20 (96GB).
Dataset (Seqlen) ELFATT FA-2 Linformer Nyström Primal Reformer VaniATT
Time (s) ||| Memory (GB)
ListOps (2K) 16.2 ||| 0.4 24.2 ||| 0.4 19.6 ||| 1.1 28.3 ||| 0.6 21.2 ||| 0.5 34.9 ||| 1.4 71.2 ||| 4.3
Text (4K) 28.8 ||| 0.8 63.7 ||| 0.8 35.1 ||| 2.2 50.4 ||| 1.2 33.9 ||| 0.9 69.3 ||| 2.8 263.2 ||| 16.5
Retrieval (4K) 55.2 ||| 1.5 126.3 ||| 1.4 69.1 ||| 4.1 96.0 ||| 2.4 66.5 ||| 1.9 137.8 ||| 5.4 523.0 ||| 17.2
Image (1K) 42.9 ||| 0.8 48.9 ||| 0.8 45.8 ||| 2.2 69.7 ||| 1.4 49.8 ||| 1.0 82.9 ||| 2.9 109.2 ||| 4.5
Pathfinder (1K) 49.5 ||| 0.8 63.8 ||| 0.8 63.4 ||| 2.2 91.0 ||| 1.4 64.6 ||| 1.0 113.5 ||| 2.9 144.2 ||| 4.5
Table 5. The comparison of object detection performance of all methods on MS COCO 2017. Note: FLOPs are calculated using an input size of 1280×8001280$\times$ 8001280 × 800. ‘1$\times$’ denotes the fine-tuning training schedule with 12 epochs and ‘3$\times$×MS’ represents fine-tuning using the multiscale training schedule with 36 epochs. APb denotes box average precision and APm denotes mask average precision. Inference throughput (FPS (nFA/FA), imgs/s) is obtained using a batch size of 1 with mixed precision on a single NVIDIA H20 (96 GB) GPU.
Mask-RCNN (He et al., 2017) 1$\times$ schedule
[…]
CSWin-T
Agent 46.8 68.9 51.3 42.3 65.9 45.3 19/20 40M 273.87G/254.51G
EFFATT 46.1 68.3 50.5 41.9 65.5 45.3 32/36 40M 329.95G/255.20G
ELFATT 47.0 69.2 51.4 42.6 66.4 45.9 30/33 40M 334.86G/254.40G
FLatten 46.6 68.8 51.0 42.2 65.7 45.3 26/28 41M 274.41G/255.05G
GLOBAL 47.0 69.1 51.9 42.6 66.1 45.9 7/18 40M 1712.76G/253.56G
LOCAL 46.5 68.5 51.0 42.1 65.6 45.3 26/28 40M 281.45G/253.57G
Swin-T
Agent 44.6 67.5 48.7 40.7 64.4 43.4 5/— 48M 278.42G/ —
EFFATT 44.7 67.0 48.9 41.1 64.0 44.4 40/46 48M 301.89G/261.95G
ELFATT 46.1 68.3 50.8 42.1 65.4 45.3 39/45 50M 311.39G/266.43G
FLatten 44.2 67.3 48.5 40.2 63.8 43.0 41/— 49M 266.43G/ —
GLOBAL 45.4 67.9 49.7 41.6 65.0 44.8 6/17 48M 2106.75G/260.48G
LOCAL 42.7 65.2 46.8 39.3 62.2 42.2 45/— 48M 267.01G/ —
Other tiny models
ConvNeXt 44.2 66.6 48.3 40.1 63.3 42.8 44/— 48M 262.29G/ —
VMamba 47.4 69.5 52.0 42.7 66.3 46.0 35/— 50M 271.16G/ —
Mask-RCNN (He et al., 2017) 3$\times$×MS schedule
[…]
CSWin-T
Agent 49.3 70.8 53.9 43.9 67.9 47.3 19/20 40M 273.87G/254.51G
EFFATT 48.5 70.0 53.2 43.4 67.3 46.9 32/36 40M 329.95G/255.20G
ELFATT 49.4 70.9 54.4 44.0 68.0 47.5 30/33 40M 334.86G/254.40G
FLatten 48.9 70.8 53.5 43.9 67.9 47.3 26/28 41M 274.41G/255.05G
GLOBAL 48.8 70.0 53.5 43.6 67.4 47.1 7/18 40M 1712.76G/253.56G
LOCAL 49.3 70.8 54.3 44.0 67.8 47.5 26/28 40M 281.45G/253.57G
Swin-T
Agent 47.3 69.5 51.9 42.7 66.4 46.2 5/— 48M 278.42G/ —
EFFATT 47.6 69.4 52.6 42.7 65.9 46.1 40/46 48M 301.89G/261.95G
ELFATT 48.5 70.4 53.4 43.6 67.3 47.3 39/45 50M 311.39G/266.43G
FLatten 46.5 68.5 50.8 42.1 65.4 45.1 41/— 49M 266.43G/ —
GLOBAL 48.0 70.0 52.7 43.3 67.0 46.8 6/17 48M 2106.75G/260.48G
LOCAL 46.0 68.1 50.3 41.6 65.1 44.9 45/— 48M 267.01G/ —
Other tiny models
ConvNeXt 46.2 67.9 50.8 41.7 65.0 44.9 44/— 48M 262.29G/ —
VMamba 48.9 70.6 53.6 43.7 67.7 46.8 35/— 50M 271.16G/ —
5.3. Performance Comparison on LRA
The block sizes of ELFATT are set to 64 as used in Nyströmformer (Nyström) in all LRA tasks to maintain fairness. As Table 3 shows, ELFATT significantly outperforms all efficient attention mechanisms in all tasks except Text (4K) and Pathfinder (1K) on LRA. All efficient attention mechanisms are inferior to VaniATT in Pathfinder (1K). ELFATT is only inferior to VaniATT in Pathfinder (1K). For Text (4K), ELFATT is slightly inferior to VaniATT. The overall performance of ELFATT is the best and significantly better than all comparison methods, including VaniATT. Table 4 shows the running time (s) per 1K training steps and peak memory consumption of all methods in the training of LRA. ELFATT is significantly faster than all comparison methods. Especially, ELFATT offers 1.2-2.3x speedups over FlashAttention-2. For most tasks, ELFATT and FlashAttention-2 consume less GPU memory than other comparison methods.
5.4. Object Detection Performance
Table 5 shows the object detection performance of all methods on MS COCO 2017. Using the Mask-RCNN (He et al., 2017) 1$\times$ schedule, VaniATT and ELFATT significantly outperform other attention mechanisms. ELFATT exhibits slightly better performance than VaniATT when using CSWin-T as the backbone and significantly outperforms VaniATT when using Swin-T as the backbone. Under the backbone of CSWin-T, ELFATT offers a 4.3x speedup over VaniATT without using FlashAttention-2 and is still 1.8x faster than VaniATT using FlashAttention-2. When using Swin-T as the backbone, ELFATT offers a 6.5x speedup over VaniATT without using FlashAttention-2 and is still 2.6x faster than VaniATT using FlashAttention-2. Using the Mask-RCNN (He et al., 2017) 3$\times$ multiscale training schedule, ELFATT still achieves the best performance. VaniATT still outperforms most linear attention mechanisms when using Swin-T as the backbone. In the 1$\times$ schedule, ELFATT using CSWin-T as the backbone achieves close performance compared to VMamba-T and significantly outperforms ConvNeXt-T. In the 3$\times$ multiscale training schedule, ELFATT using CSWin-T as the backbone outperforms ConvNeXt-T and VMamba-T in terms of object detection performance.
Table 6. The comparison of inference throughput of different attention mechanisms obtained on ImageNet-1K.
Method Res. FPS (FP32) Speedup FPS (Mixed) Speedup
Backbone: CSWin-T; Platform: NVIDIA Jetson AGX Orin; Batch Size: 128; Mode: 60W
Agent $224^{2}$ 158 imgs/s 1.5x 295 imgs/s 1.3x
EFFATT $224^{2}$ 156 imgs/s 1.4x 305 imgs/s 1.3x
ELFATT $224^{2}$ 174 imgs/s 1.6x 355 imgs/s 1.6x
FLatten $224^{2}$ 138 imgs/s 1.3x 229 imgs/s 1.0x
GLOBAL $224^{2}$ 108 imgs/s 1.0x 298 imgs/s 1.3x
LOCAL $224^{2}$ 167 imgs/s 1.5x 309 imgs/s 1.3x
Backbone: Others; Platform: NVIDIA Jetson AGX Orin; Batch Size: 128; Mode: 60W
EfficientViT-B2 $288^{2}$ 198 imgs/s 1.8x 282 imgs/s 1.2x
Backbone: CSWin-T; Platform: NVIDIA Jetson Nano; Batch Size: 1; Mode: 10W
Agent $224^{2}$ 4.9 imgs/s 1.5x 4.5 imgs/s 1.3x
EFFATT $224^{2}$ 5.8 imgs/s 1.8x 5.5 imgs/s 1.6x
ELFATT $224^{2}$ 6.2 imgs/s 2.0x 5.8 imgs/s 1.7x
FLatten $224^{2}$ 4.7 imgs/s 1.5x 4.0 imgs/s 1.1x
GLOBAL $224^{2}$ 3.2 imgs/s 1.0x 3.5 imgs/s 1.0x
LOCAL $224^{2}$ 5.0 imgs/s 1.6x 4.2 imgs/s 1.2x
Backbone: Others; Platform: NVIDIA Jetson Nano; Batch Size: 1; Mode: 10W
EfficientViT-B2 $288^{2}$ 5.5 imgs/s 1.7x 5.7 imgs/s 1.6x
Table 7. Inference speed on ImageNet-1K using 1 Jetson AGX Orin (60W) with a batch size of 128 in mixed precision using TensorRT to speed up CSWin-T-ELFATT.
Running speed TensorRT Pytorch
Running Time (s/batch) 0.23 0.36
Inference Throughput (imgs/s) 558 352
Total Inference Time (s) 90 142
5.5. Speed Comparison on Edge GPUs
Table 6 shows the comparison of inference speed of different attention mechanisms obtained on ImageNet-1K using an NVIDIA Jetson AGX Orin/NVIDIA Jetson Nano GPU. In both FP32 and mixed precision, under the backbone of CSWin-T, ELFATT achieves the highest FPS. In mixed precision, ELFATT is significantly faster than all other attention mechanisms, as shown in Table 6. Even compared to edge-optimized EfficientViT-B2 (Cai et al., 2023), to achieve similar accuracy, ELFATT is significantly faster than EfficientViT-B2 in mixed precision on edge GPUs. Fig. 5 in Appendix shows the speed comparison of different attention mechanisms across various power modes from 5W to 60W by using NVIDIA Jetson Nano or NVIDIA Jetson AGX Orin. ELFATT consistently achieves 1.6x to 2.0x speedups compared to other attention mechanisms in various power modes from 5W to 60W. Compared to edge-optimized EfficientViT-B2, in mixed precision, ELFATT performs on par with EfficientViT-B2 in low-power conditions, such as the 5W mode. As power increases, the speed growth rate of ELFATT surpasses that of EfficientViT-B2, especially in high-power modes like 50W and 60W, where it demonstrates a significant speed advantage. Furthermore, in mixed precision, ELFATT outperforms all other attention mechanisms in terms of speed across all power modes tested, establishing itself as one of the fastest options in both high-performance and power-constrained scenarios. Table 7 shows that ELFATT can be further accelerated to achieve a 1.6x speedup by using TensorRT on a Jetson AGX Orin.
Table 8. The comparison of stable diffusion (SD) v1.5 (Diffusers v0.32.1) using different acceleration methods with different merging ratios in FP16 precision. Memory consumption for each method indicates how much GPU memory increases when the batch size is raised by 1. All methods are accelerated by FlashAttention-2. Note: The processing time of each method is obtained using a batch size of 48 on 1 NVIDIA Tesla A100 (40 GB) GPU.
Method Ratio FID $\downarrow$ Time (s/img) Memory (GB/img)
Agent-SD 0.1 30.2 0.794 1.00
Agent-SD 0.2 30.2 0.777 1.00
Agent-SD 0.3 30.3 0.760 1.00
Agent-SD 0.4 30.6 0.746 0.70
Agent-SD 0.5 30.7 0.724 0.70
ELFATT-SD 0.1 30.2 0.772 0.80
ELFATT-SD 0.2 30.2 0.756 0.75
ELFATT-SD 0.3 30.4 0.739 0.75
ELFATT-SD 0.4 30.2 0.725 0.70
ELFATT-SD 0.5 30.4 0.707 0.70
SD (Baseline) 31.0 0.768 0.80
ToMe-SD 0.1 30.7 0.788 0.80
ToMe-SD 0.2 30.5 0.765 0.80
ToMe-SD 0.3 30.6 0.748 0.80
ToMe-SD 0.4 30.6 0.732 0.80
ToMe-SD 0.5 30.9 0.714 0.70
5.6. Stable Diffusion Acceleration Using ELFATT
Table 8 shows the performance comparison of stable diffusion (SD) (Rombach et al., 2022) v1.5 (Diffusers v0.32.1) by using different acceleration methods. Compared to Agent (Agent-SD) and ToMe (ToMe-SD) (Bolya and Hoffman, 2023), ELFATT (ELFATT-SD) provides lower FID (Heusel et al., 2017) scores, consumes less GPU memory, and shows faster computation speed in the same merging ratio. From Table 8, it can be seen that SD using FlashAttention-2 is able to achieve a comparable speed and memory usage compared to ToMe and Agent when the merging ratio is small. Only ELFATT is faster and has less memory usage than SD using FlashAttention-2 in most merging ratios. Fig. 6 in Appendix shows the visual comparison of images generated by different methods. ELFATT can help SD generate richer details and reduce unreality. Taking the prompt “Komodo dragon” as an example, ELFATT helps SD generate a “Komodo dragon” with a more realistic body proportion, which is more like a real monitor.
6. Conclusion
A novel efficient linear fast attention (ELFATT) mechanism is proposed for ViTs to achieve low memory I/O operations, linear computational complexity, and high performance at the same time. ELFATT offers 4-7x speedups over VaniATT in high-resolution vision tasks without losing performance. ELFATT is compatible with FlashAttention. Using FlashAttention-2, ELFATT still offers 2-3x speedups over VaniATT in high-resolution vision tasks without losing performance. ELFATT without using FlashAttention-2 is even faster than VaniATT using FlashAttention-2 on both low-resolution and high-resolution vision datasets, which shows great potential of ELFATT. Even in some non-vision tasks of LRA, ELFATT still achieves leading performance and offers 1.2-2.3x speedups over FlashAttention-2. What’s more, on edge GPUs, ELFATT still offers 1.6x to 2.0x speedups compared to state-of-the-art attention mechanisms in various power modes from 5W to 60W. In addition, ELFATT can be used to enhance and accelerate diffusion tasks directly without training.
Acknowledgement
This work is supported by the Hong Kong Innovation and Technology Commission (InnoHK Project CIMDA), the Institute of Digital Medicine, City University of Hong Kong (Project 9229503), and the Natural Science Special Project Research Fund of Guizhou University (No. 2025-06).
Appendix A Approximation Error Bound Analysis
Eq. (8) of the main body of the paper can be derived as follows,
$\begin{split}&{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}=\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\textit{{V}}\\ &={\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\left[\bar{\textit{{V}}},\tilde{\textit{{V}}}\right]\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\quad\left({\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\\ &\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right],\\ \end{split}$
where $\odot$ denotes the Hadamard product (Horn and Johnson, 2012), ${\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\tilde{\textit{{V}}}$ is equivalent to $g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)$, and $\textit{{Z}}\in\mathbb{R}^{m\times m}$ is a matrix as follows,
$\textit{{Z}}=\textbf{I}_{b}\otimes\textit{{U}}_{(m/b)},$
with $\otimes$ denoting the Kronecker product (Horn and Johnson, 2012) and $\textit{{U}}_{(m/b)}\in\mathbb{R}^{(m/b)\times(m/b)}$ being the all-ones matrix. It is obvious to see that
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\right.\\ &\quad\left.\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right],\end{split}$
which implies that
(12) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi},\end{split}$
where $\xi=2$ denotes the spectral norm and $\xi=F$ denotes the Frobenius norm. For given two matrices $\textit{{A}}\in\mathbb{R}^{m\times n}$ and $\textit{{B}}\in\mathbb{R}^{n\times p}$, it follows from (Golub and Van Loan, 2013) [Section 2.3.1] that $\|\textit{{A}}\textit{{B}}\|_{\xi}\leq\|\textit{{A}}\|_{\xi}\|\textit{{B}}\|_{\xi}$.
For the second term in the right-hand side of Inequality (12), it is easy to see that
$\begin{split}&\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
For the first term in the right-hand side of Inequality (12), we have the following theorem.
Theorem A.1.
Let $\emph{{{U}}}_{m}\in\mathbb{R}^{m\times m}$ be an all-ones matrix. For any two vectors $\bar{\textit{{q}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$ and $\bar{\textit{{k}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, let $\mathfrak{M}>0$ and $\mathscr{M}>0$ be the maximum and minimum of ${\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}+0.5-\left(\bar{q}_{i}+\bar{k}_{i}\right)\right)$, respectively, and $\bar{q}_{i}$ and $\bar{k}_{i}$ are the elements at position $i$ of vectors $\bar{\textit{{q}}}$ and $\bar{\textit{{k}}}$, respectively. If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the following inequality holds,
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
Before proving Theorem A.1, we introduce the following lemma (see (Wu et al., 2024)).
Lemma A.2.
For any two vectors $\bar{\textit{{q}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$ and $\bar{\textit{{k}}}\in\mathbb{R}^{c_{1}}$ from $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, let
$\begin{split}D_{\bar{\textit{{q}}},\bar{\textit{{k}}}}&=\max_{i=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}+0.5-\left(\bar{q}_{i}+\bar{k}_{i}\right)\right)>0,\\ d_{\bar{\textit{{q}}},\bar{\textit{{k}}}}&=\min_{i=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}+0.5-\left(\bar{q}_{i}+\bar{k}_{i}\right)\right)>0,\end{split}$
where $\bar{q}_{i}$ and $\bar{k}_{i}$ are the elements at position $i$ of vectors $\bar{\textit{{q}}}$ and $\bar{\textit{{k}}}$, respectively. The following inequalities hold,
$\begin{split}\frac{{\rm exp}\left(\bar{\textit{{q}}}\right){\rm exp}\left(\bar{\textit{{k}}}\right)^{\top}}{{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}\right)}&\leq\frac{c_{1}}{d_{\bar{\textit{{q}}},\bar{\textit{{k}}}}{\rm exp}\left(-0.5\right)},\\ \frac{{\rm exp}\left(\bar{\textit{{q}}}\right){\rm exp}\left(\bar{\textit{{k}}}\right)^{\top}}{{\rm exp}\left(\bar{\textit{{q}}}\bar{\textit{{k}}}^{\top}\right)}&\geq\frac{c_{1}}{D_{\bar{\textit{{q}}},\bar{\textit{{k}}}}{\rm exp}\left(-0.5\right)}.\end{split}$
The following corollary is easily obtained from Lemma A.2.
Corollary A.3.
For two matrices $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$ and $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, let
$\begin{split}\mathfrak{M}&=\max_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}_{i_{1}}\bar{\textit{{k}}}_{i_{2}}^{\top}+0.5-\left(\bar{q}_{i_{1}j}+\bar{k}_{i_{2}j}\right)\right)>0,\\ \mathscr{M}&=\min_{i_{1},i_{2}=1,2,\dots,m\atop j=1,2,\dots,c_{1}}{\rm exp}\left(\bar{\textit{{q}}}_{i_{1}}\bar{\textit{{k}}}_{i_{2}}^{\top}+0.5-\left(\bar{q}_{i_{1}j}+\bar{k}_{i_{2}j}\right)\right)>0,\end{split}$
where $\bar{q}_{i_{1}j}$ and $\bar{k}_{i_{2}j}$ are the elements at position $j$ of the vector $\bar{\textit{{q}}}_{i_{1}}=\bar{\textit{{Q}}}(i_{1},:)$ and the vector $\bar{\textit{{k}}}_{i_{2}}=\bar{\textit{{K}}}(i_{2},:)$, respectively. The following inequalities hold,
$\begin{split}\frac{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}}&\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)},\\ \frac{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\right\|_{\xi}}{\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}}&\geq\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}.\end{split}$
Proof.
For any matrices $\bar{\textit{{Q}}}\in\mathbb{R}^{m\times c_{1}}$, $\tilde{\textit{{Q}}}\in\mathbb{R}^{m\times c_{2}}$, $\bar{\textit{{K}}}\in\mathbb{R}^{m\times c_{1}}$, and $\tilde{\textit{{K}}}\in\mathbb{R}^{m\times c_{2}}$, we have
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &=\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)+\right.\\ &\left.{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\\ &\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\right\|_{\xi}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
Following from Corollary A.3, if $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &\leq\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\emph{{{U}}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\\ &+\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}.\\ \end{split}$
The proof is completed. ∎
We now consider the upper bound for the right-hand side of Inequality (12). If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,
(13) $\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\\ &\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,
(14) $\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\\ &\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
For the approximation error of Eq. (9) of the main body of the paper, it is obvious to see that
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right].\end{split}$
Hence, the corresponding approximation error is bounded as follows,
(15) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
(16) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\\ &\leq\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-\textit{{U}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
(17) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\\ &\leq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-\textit{{U}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
Inequalities (15-17) give a tighter bound than Inequalities (12-14). Fig. 2 (a) shows the upper bounds of different attention mechanisms (EFFATT (Shen et al., 2021), ELFATT, and LOCAL (Dong et al., 2022)) for approximating attention matrices of vanilla attention during the training of ImageNet-1K. The backbone used is CSWin-T (Han et al., 2023). The upper bound for ELFATT is obtained according to
$\begin{split}&\left\lVert\left[\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top},\ \left(\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}\right)\right]-\left[\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}),\ \mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right]\right\rVert_{\xi}\\ &\leq\left\lVert\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top}-\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\right\rVert_{\xi}+\left\lVert\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}-\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right\rVert_{\xi},\end{split}$
for EFFATT is as
$\begin{split}&\left\lVert\left[\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top},\ \mathrm{exp}(\tilde{\textit{{Q}}})\mathrm{exp}(\tilde{\textit{{K}}})^{\top}\right]-\left[\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}),\ \mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right]\right\rVert_{\xi}\\ &\leq\left\lVert\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top}-\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\right\rVert_{\xi}+\left\lVert\mathrm{exp}(\tilde{\textit{{Q}}})\mathrm{exp}(\tilde{\textit{{K}}})^{\top}-\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right\rVert_{\xi},\end{split}$
and for LOCAL is as
$\begin{split}&\left\lVert\left[\left(\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\odot\textit{{Z}}_{1}\right),\ \left(\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}_{2}\right)\right]-\left[\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}),\ \mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right]\right\rVert_{\xi}\\ &\leq\left\lVert\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\odot\textit{{Z}}_{1}-\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\right\rVert_{\xi}+\left\lVert\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}_{2}-\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right\rVert_{\xi},\end{split}$
where $\xi=F$ denotes the Frobenius norm, and $\textit{{Z}}_{1}$ and $\textit{{Z}}_{2}$ are different to Z of ELFATT, because LOCAL uses a more complex cross-shaped blockify method for different heads. All upper bounds were obtained from the final attention layer of the second level of the backbone. The upper bounds of different methods show a decreasing trend as the training process progresses. ELFATT further reduces the upper bound of the LOCAL attention mechanism, although it still has a higher bound than EFFATT, which is caused by its sparse blockify attention heads. Figs. 2 (b) and (c) show the comparison of relative attention matrix approximation error ($\left\lVert\textit{{A}}-\textit{{A}}^{{}^{\prime}}\right\rVert_{\xi}/\left\lVert\textit{{A}}\right\rVert_{\xi}$, where A denotes the attention matrix of VaniATT and $\textit{{A}}^{\prime}$ denotes the attention matrix of linear attention) of global linear attention heads of ELFATT and the corresponding heads in EFFATT and relative attention matrix approximation error of sparse blockify attention heads of ELFATT and the corresponding heads in LOCAL. The backbone used is CSWin-T (Han et al., 2023). Since ELFATT uses a hybrid head architecture, its half heads consist of global linear attention heads, and the remaining heads consist of sparse blockify attention heads. We compared the relative attention matrix approximation error of global linear attention heads of ELFATT and the heads in the same position of EFFATT. We compared the relative attention matrix approximation error of sparse blockify attention heads of ELFATT and the heads in the same position of LOCAL. We found that the global linear attention heads of ELFATT have a lower approximation error than the heads in the same position of EFFATT, and the sparse blockify attention heads of ELFATT have a lower approximation error than the heads in the same position of LOCAL. The hybrid head architecture has complementary effects on approximation error reduction.
(a) Upper Bounds
(b) Global Linear Attention Heads
(c) Sparse Blockify Attention Heads
Figure 2. (a) The upper bound of different attention mechanisms (EFFATT (Shen et al., 2021), ELFATT, and LOCAL (Dong et al., 2022)) for approximating attention matrices of vanilla attention during the training of ImageNet-1K. (b) The comparison of relative attention matrix approximation error of global linear attention heads of ELFATT and the corresponding heads in EFFATT. (c) The comparison of relative attention matrix approximation error of sparse blockify attention heads of ELFATT and the corresponding heads in LOCAL. The backbone used is CSWin-T.
(a) True Label: American Staffordshire Terrier
(b) True Label: Hartebeest
(c) True Label: Zebra
(d) True Label: Library
(e) True Label: Snail
(f) True Label: Black-Footed Ferret
(g) True Label: Scottish Deerhound
(h) True Label: Pineapple
Figure 3. Visual comparison of class activation map (CAM) based attention results of different attention mechanisms obtained using Score-CAM (Wang et al., 2020b). Note: Backbone used is CSWin-T (Han et al., 2023).
(a) CSWin
(b) Swin
Figure 4. Accuracy-efficiency curves obtained by different attention mechanisms on ImageNet-1K using different bacbones (CSWin (Han et al., 2023) and Swin (Liu et al., 2021)). Efficiency is represented by the inference throughput of each method. All methods’ inference throughput and accuracy results are from Table 1 of the main body of the paper.
Appendix B Ablation Study of Different Block Sizes
To validate the effect of the block size on the performance of ELFATT, we performed an ablation study using different block sizes at each level of CSWin-T-ELFATT on ImageNet-1K. As shown in Table 9, with the increasing number of block sizes, the inference speed of ELFATT will slow down, while the classification accuracy of ELFATT shows an increasing trend. ELFATT using 49-49-196-49 and 196-196-196-49 blocks achieve the best performance in terms of speed and accuracy. It seems that FlashAttention-2 is more efficient for some block sizes (with FlashAttention-2, the block size of 196 is faster than the block size of 49 which may be caused by GPU architectures). Hence, the block sizes for ELFATT can be determined according to the performance and efficiency requirements.
Appendix C Ablation Study of LePE
To validate the effect of LePE on performance, we compared the performance of ELFATT and VaniATT using or without using LePE in CSWin-T on ImageNet-1K. As shown in Table 10, using LePE, the performance of VaniATT improves from 82.9 to 83.1 and the performance of ELFATT also improves from 82.9 to 83.1, respectively, further demonstrating the powerful local positional information enhancement of LePE. More details about LePE can be found in (Dong et al., 2022).
Table 9. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT using different block sizes in each level (L1-L4) of CSWin-T on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold.
Block size of each level Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
L1 L2 L3 L4
49 49 49 49 $224^{2}$ 82.5 2619/2805 imgs/s 20M 4.37G/4.13G
49 49 196 49 $224^{2}$ 83.1 2603/2856 imgs/s 20M 4.44G/4.13G
49 196 196 49 $224^{2}$ 82.8 2552/2863 imgs/s 20M 4.50G/4.13G
196 196 196 49 $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
Table 10. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT and VaniATT using or without using LePE in CSWin-T on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “w/o” denotes “without”, “nFA” denotes non-FlashAttention-2, and “FA” denotes FlashAttention-2.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
CSWin-T-ELFATT
LePE $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
w/o-LePE $224^{2}$ 82.9 2801/3271 imgs/s 20M 4.54G/4.12G
CSWin-T-GLOBAL
LePE $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
w/o-LePE $224^{2}$ 82.9 1332/2296 imgs/s 20M 7.58G/4.08G
Table 11. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by using the ELFATT modules to replace the VaniATT modules in different levels (L1-L4) of CSWin-T-GLOBAL on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold. “Res.” denotes resolution, “imgs” denotes images, “nFA” denotes non-FlashAttention-2, and “FA” denotes FlashAttention-2. ✓denotes the level consists of full ELFATT modules and ✔ denotes half of this level is composed of ELFATT modules and the remaining is composed of VaniATT modules. The order of composition is: VaniATT-ELFATT-...-VaniATT-ELFATT.
Levels using ELFATT Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
L1 L2 L3 L4
$224^{2}$ 83.0 2257/2865 imgs/s 20M 5.17G/4.10G
$224^{2}$ 83.0 2467/2918 imgs/s 20M 4.63G/4.12G
$224^{2}$ 82.6 2555/2842 imgs/s 20M 4.48G/4.15G
$224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
$224^{2}$ 82.6 2552/2835 imgs/s 20M 4.48G/4.15G
CSWin-T-GLOBAL $224^{2}$ 83.1 1303/2210 imgs/s 20M 7.60G/4.09G
Table 12. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) obtained by CSWin-T-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on a single NVIDIA H20 (96 GB) GPU. The best values are in bold.
Combinations of $c_{1}$ and $c_{2}$ Res. Acc. FPS (nFA/FA) # FLOPs (nFA/FA)
$c_{1}$ $c_{2}$
$0\times c$ $1\times c$ $224^{2}$ 82.7 2331/2880 imgs/s 20M 4.76G/4.09G
0.25×0.25$\times$0.25 × c $0.75\times c$ $224^{2}$ 83.0 2530/2835 imgs/s 20M 4.65G/4.10G
0.5×0.5$\times$0.5 × c $0.5\times c$ $224^{2}$ 83.1 2512/2881 imgs/s 20M 4.56G/4.13G
0.75×0.75$\times$0.75 × c $0.25\times c$ $224^{2}$ 83.1 2458/2796 imgs/s 20M 4.48G/4.18G
1×1$\times$1 × c $0\times c$ $224^{2}$ 82.6 2394/2526 imgs/s 20M 4.35G/4.17G
Appendix D Ablation Study of Number of Levels Using the ELFATT Module
To validate the effect of the number of ELFATT modules used to replace the VaniATT modules at different levels of vision transformers, we compared the performance of CSWin-T using a different number of ELFATT modules at different levels. As shown in Table 11, with the number of levels using ELFATT modules increasing, the inference speed increases without using FlashAttention-2. The difference between CSWin-T using 3 levels of ELFATT modules and 4 levels of ELFATT modules is not significant. The fourth level of CSWin-T is composed of only one module and the sequence length is only 49 which is too short to achieve a significant acceleration effect. Another thing that can be found in Table 11 is that with an increase in the number of levels using ELFATT modules, CSWin-T using FlashAttention-2 achieves the fastest inference speed when using two levels of ELFATT modules. Because the sequence lengths of the first two levels are 3136 and 784, respectively, which are significantly longer than the last two levels (the $3^{\rm rd}$ level: 196, and the $4^{\rm th}$ level: 49). That is also the reason why some efficient attention mechanisms, such as Agent (Han et al., 2025) and FLatten (Han et al., 2023), only replace some of levels by their efficient attention modules. We also observed that with an increasing number of levels using ELFATT modules, the performance shows a gradual decline. To address this defect, we introduced a hybrid architecture in the third level which replaces half of the ELFATT modules at this level by VaniATT modules. The order of composition is as follows: VaniATT-ELFATT-...-VaniATT-ELFATT. The reason is that the third level of the vision transformer is usually much deeper than other levels and the sequence of this level is also much shorter than the first two levels. VaniATT at this level will not affect the speed too much and can help the model to converge faster. Swin-T-ELFATT, Swin-B-ELFATT, and CSWin-B-ELFATT also use a pure ELFATT architecture in the first two levels and a hybrid architecture in the third level. All variants of CSWin-T using ELFATT modules to replace VaniATT modules are faster than CSWin-T-GLOBAL.
Appendix E Ablation Study of Different $c_{1}$ and $c_{2}$
Table 12 shows the performance comparison of CSWin-T-ELFATT using different combinations of $c_{1}$ and $c_{2}$ on ImageNet-1K. When $c_{1}=1\times c$ and $c_{2}=0\times c$, ELFATT becomes EFFATT, and when $c_{1}=0\times c$ and $c_{2}=1\times c$, ELFATT becomes a local window-based attention mechanism which can be regarded as a simpler version of the local window-based attention mechanism used in Swin. As shown in Table 12, $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ achieve better performance than other combinations of $c_{1}$ and $c_{2}$. Without using FlashAttention-2, the difference between $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$, and $c_{1}=0.75\times c$ and $c_{2}=0.25\times c$ in terms of speed is not significant. Using FlashAttention-2, $c_{1}=0\times c$ and $c_{2}=1\times c$, and $c_{1}=0.5\times c$ and $c_{2}=0.5\times c$ achieve close speed and are significantly faster than other combinations.
Appendix F Ablation Study of Different Patch Partition/Merging Methods
Table 13 shows the performance comparison of Swin-T-ELFATT using different patch partition/merging methods on ImageNet-1K. “concat” (Liu et al., 2021) denotes the concatenation-based patch partition/merging method of Swin. “conv” (Dong et al., 2022) denotes the convolution-based patch partition/merging method of CSWin. It can be seen that “conv” significantly outperforms “concat” in terms of speed and accuracy. Hence, in this paper, all ELFATT variants use “conv” for patch partition/merging.
Appendix G Speed Comparison on Edge GPUs: Experiment Settings
In addition to high-performance computing applications, model deployment should also be evaluated in emerging edge scenarios such as robotic vision, unmanned aerial vehicles (UAVs), and autonomous driving. These scenarios typically operate under strict power constraints while striving to maximize performance and efficiency. The NVIDIA Jetson series provide a good embedded and edge computing platform to evaluate AI model performance in such resource-constrained environments.
Table 13. The comparison of top-1 test accuracy (Acc.), inference throughput (FPS), parameter numbers (#), and number of floating point operations (FLOPs) of ELFATT using different patch partition/merging methods in Swin-T on ImageNet-1K. Note: Inference throughput is obtained using a batch size of 512 with mixed precision on 1 NVIDIA H20 (96 GB) GPU.
Method Res. Acc. (%) FPS (nFA/FA) # FLOPs (nFA/FA)
concat $224^{2}$ 82.3 2027/2769 imgs/s 28M 6.81G/4.42G
conv $224^{2}$ 82.7 2884/3159 imgs/s 30M 4.99G/4.67G
G.1. Experiment Settings
The experiments were conducted to evaluate the inference speed of the model on NVIDIA Jetson platforms, specifically Jetson Nano and Jetson AGX Orin, across power modes ranging from 5W to 60W. Each model was evaluated in full precision (FP32) and mixed precision modes. The evaluations were performed on the ImageNet-1K dataset with batch sizes of 1 for Jetson Nano and 128 for Jetson AGX Orin. Both devices installed the latest NVIDIA JetPack SDK (The Jetson Nano used JetPack SDK 4.6.6 and the Jetson AGX Orin utilized JetPack SDK 6.1). Each experiment was repeated 100 times and the reported results represent the average values to ensure statistical reliability.
Figure 5. The speed comparison of different attention mechanisms across power modes from 5W to 60W on NVIDIA Jetson Nano and AGX Orin under power-constrained inference scenarios.
Appendix H Stable Diffusion Acceleration Using ELFATT: Experiment Settings and Ablation Studies
We also introduced ELFATT to the acceleration of stable diffusion (SD) (Rombach et al., 2022). The original EFFATT-based global linear attention head and LePE of ELFATT cannot be used without training. We used Agent (Han et al., 2025) to replace the original EFFATT-based global linear attention head. The reason is that: (1) Agent is an improved version of EFFATT, and it only introduces a small agent matrix as the auxiliary key or query matrix to be multiplied with the full query or full key matrix to reduce its dimenison before the softmax normalization; (2) Agent can be used to replace the self-attention block of SD without training. Hence, Agent can be introduced to replace the original EFFATT-based global linear attention head of ELFATT. LePE is removed. Finally, in the self-attention block of SD, Eq. (10) of the main body of the paper will become as follows,
(18) $\begin{split}{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}\approx&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\mathcal{A}}^{\top}\right){\rm exp}\left(\bar{\mathcal{A}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$
Eq. (11) of the main body of the paper will become as follows,
(19) $\begin{split}\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\approx&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\mathcal{A}}^{\top}\right){\rm exp}\left(\bar{\mathcal{A}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right],\end{split}$
where $\bar{\mathcal{A}}={\rm Pooling}\left(\bar{\textit{{Q}}}\right)\in\mathbb{R}^{m_{1}\times c_{1}}$, ${\rm Pooling}(\cdot)$ denotes the average pooling function, and $m_{1}\ll m$.
Table 14. The comparison of stable diffusion (SD) v1.5 (Diffusers v0.32.1) using ELFATT with different block sizes in FP16 precision where merging ratio is 0.5. Memory consumption for each variant indicates how much GPU memory increases when the batch size is raised by 1. All variants are accelerated by FlashAttention-2. The processing time of each variant is obtained using a batch size of 48 on 1 NVIDIA Tesla A100 (40 GB) GPU.
Block size FID $\downarrow$ Time (s/img) Memory (GB/img)
32 30.8 0.710 0.65
56 30.7 0.709 0.65
81 30.5 0.708 0.65
128 30.4 0.707 0.70
227 30.5 0.705 0.70
512 30.5 0.709 0.70
H.1. Experiment Settings
The experiments were carried out to assess the performance and acceleration effect of ELFATT in a stable diffusion task using 1 NVIDIA Tesla A100 (40 GB) GPU. Following the experiment settings of (Han et al., 2025) and (Bolya and Hoffman, 2023), we applied ELFATT (Eqs. (18) and (19)) on stable diffusion v1.5 using the pipeline of Diffusers v0.32.1 and called the proposed stable diffusion acceleration method ELFATT-SD. We compared ELFATT with Agent (Agent-SD) (Han et al., 2025) and ToMe (ToMe-SD) (Bolya and Hoffman, 2023). We used ImageNet-1K labels as prompts to generate 2000 images with a resolution of $512^{2}$ (2 images per class). The inference steps (50) (Liu et al., 2022b) and the cfg scale (7.5) (Dhariwal and Nichol, 2021) remain the same as Agent and ToMe. Similar to Agent-SD, ELFATT-SD is also developed based on the token merging method ToMe-SD. We also used the hybrid architecture of Agent-SD which means ELFATT was used in early inference steps, and after that vanilla ToMe was applied for the remaining inference steps. The ratio of inference steps used for ELFATT is 20%. The evaluation metric used is the FID score (Heusel et al., 2017) which is used to evaluate the difference between 2000 generated images and 50000 images of the ImageNet-1K validation set.
Figure 6. The visual comparison of images generated by Agent-SD (merging ratio: 0.5), ELFATT-SD (merging ratio: 0.5), SD, and ToMe-SD (merging ratio: 0.5).
Table 15. The comparison of SD using ELFATT in different inference steps with BF16 precision (merging ratio is 0.5).
Steps FID $\downarrow$ Time (s/img) Memory (GB/img)
0 (ToMe-SD) 30.9 0.714 0.70
10 30.4 0.707 0.70
20 30.6 0.702 0.60
H.2. Ablation Study of Different Block Sizes
Table 14 shows the effect of different block sizes used in ELFATT on the performance of the stable diffusion task. When the block size increases, the speed is first improved and then declined. When the block size is 227, the speed is the fastest. The speed of the block size 128 is the second fastest and this block size has the lowest FID score. It seems that when the block size becomes smaller than 227, the inference speed is slowed, which may be caused by FlashAttention-2 and GPU architectures. In summary, the speed difference of the block size range from 32 to 512 is not significant. The reason is that: (1) Even the largest block size 512 is 8x smaller than the original sequence length 4096. (2) Only the first 20% inference steps were applied to use ELFATT for acceleration. However, our method is still significantly faster than other methods, especially the baseline which shows the effectiveness of ELFATT for acceleration of SD.
H.3. Ablation Study of Number of Inference Steps Using ELFATT
Table 15 shows the effect of different inference steps using ELFATT on the performance of the stable diffusion task. When 0 inference steps using ELFATT are adopted, ELFATT-SD becomes ToMe-SD. With more inference steps using ELFATT being adopted, the speed becomes faster; however, the FID score is reduced first and then increased. When early 10 (20%) inference steps using ELFATT are adopted, the FID score is the lowest.